### v1

# UFO3: Weaving the Digital Agent Galaxy [Uncaptioned image]

Abstract
Large language model (LLM)-powered agents are transforming digital devices from passive tools into proactive intelligent collaborators. However, most existing frameworks remain confined to a single OS or device, making cross-device workflows brittle and largely manual. We present UFO3 [Uncaptioned image], a system that unifies heterogeneous endpoints, desktops, servers, mobile devices, and edge, into a single orchestration fabric. UFO3 models each user request as a mutable TaskConstellation: a distributed DAG of atomic subtasks (TaskStars) with explicit control and data dependencies (TaskStarLines). The TaskConstellation continuously evolves as results stream in from distributed devices, enabling asynchronous execution, adaptive recovery, and dynamic optimization. A Constellation Orchestrator executes tasks safely and asynchronously while applying dynamic DAG updates, and the Agent Interaction Protocol (AIP) provides persistent, low-latency channels for reliable task dispatch and result streaming. These designs dissolve the traditional boundaries between devices and platforms, allowing agents to collaborate seamlessly and amplify their collective intelligence.
We evaluate UFO3 on NebulaBench, a benchmark of 55 cross-device tasks across 5 machines and 10 categories. UFO3 achieves 83.3% subtask completion, 70.9% task success, exposes parallelism with an average width of 1.72, and reduces end-to-end latency by 31% relative to a sequential baseline. Fault-injection experiments demonstrate graceful degradation and recovery under transient and permanent agent failures. These results show that UFO3 achieves accurate, efficient, and resilient task orchestration across heterogeneous devices, uniting isolated agents into a coherent, adaptive computing fabric that extends across the landscape of ubiquitous computing.
We developed UFO3 as a fully engineered system with over 73K lines of code, encompassing agent implementations and integrations for Windows, Linux, and Android mobile devices. The entire project is open-sourced at https://github.com/microsoft/UFO/, accompanied by detailed documentation and tutorials at https://microsoft.github.io/UFO/
Figure 1: UFO3: Weaving the Digital Agent Galaxy. A single natural-language intent is decomposed into a dynamically evolved Constellation (DAG) executed across heterogeneous devices. Demo video available at: https://www.youtube.com/watch?v=NGrVWGcJL8o.
1 Introduction
The rise of intelligent agents wang2024survey marks a new era of human–computer interaction, where large language models (LLMs) naveed2025comprehensive are evolving from text-based reasoning engines qu2025tool into autonomous digital operators capable of perceiving, acting, and coordinating across tasks zhang2024large. Yet despite this progress, most agent frameworks remain confined within a single device or platform, be it a browser tab ning2025survey; zheng2024gpt, a desktop environment zhang2025ufo; zhang2025ufo2, or a mobile app zhang2025appagent; wang2024mobile. This confinement sharply limits their ability to harness the rich, distributed computational ecosystem that modern users inhabit.
When an agent is trapped within one operating system, it cannot access the complementary strengths of other devices, such as GPU clusters for computation, desktop applications for document editing, or mobile sensors for context capture. The result is a fragmented landscape of intelligent but siloed agents: each powerful in isolation yet collectively underutilized. This gap between reasoning ability and real-world actuation leaves vast potential untapped. To truly advance the next frontier of automation and reasoning, agents must operate beyond the boundaries of any single device or OS, forming a coherent digital collective where Windows laptops, Linux servers, mobile devices, and edge nodes collaborate seamlessly houben2017opportunities; brudy2019cross to gather and act upon ubiquitous intelligence.
Imagine a future where you could simply say: “Prepare a production-ready demo of Project X and deliver a one-page executive summary with screenshots and performance numbers.” Today, this requires tedious, error-prone coordination across devices, checking out code on a laptop, triggering GPU builds on a server, deploying to a cloud instance, recording UI interactions on a phone, and stitching results into a report. Despite recent advances in intelligent agents, most systems remain confined within a single device or platform, leaving vast computational resources underutilized.
To realize this vision of seamless cross-device collaboration, we must overcome three interlocking challenges that go beyond classical workflow engines or single-machine agents. First, asynchronous parallelism: many subtasks can and should run concurrently across devices with varying capabilities. Second, distributed coordination: agents need reliable, low-latency communication for task dispatch and result streaming despite network variability. Third, heterogeneous extensibility: the system should make it easy to develop and integrate new device agents while preserving safety and global consistency.
We present UFO3: Weaving the Digital Agent Galaxy, a cross-device orchestration system that turns isolated devices, desktops, servers, mobile, and edge, into a coherent execution fabric. UFO3 models each request as a TaskConstellation: a dynamic distributed DAG whose nodes (TaskStars) represent executable subtasks and whose edges (TaskStarLines) capture data and control dependencies. The Constellation serves as both the logical plan and live runtime substrate: nodes are assigned asynchronously, executed opportunistically, and continuously updated as results stream. Figure 1 illustrates this concept, where one intent decomposed into a distributed DAG and orchestrated across heterogeneous endpoints.
To realize these capabilities and address the challenges in cross-device agents, UFO3 is built around five tightly integrated design principles:
  • Declarative decomposition into a dynamic DAG (TaskConstellation). Natural-language or programmatic requests are decomposed by the global ConstellationAgent into a structured DAG bei2025graphs of TaskStars and TaskStarLines that encode workflow logic and dependencies. This declarative structure is amenable to automated scheduling, introspection, and rewriting throughout execution.
  • Continuous, result-driven graph evolution. The TaskConstellation is a living data structure. Intermediate outputs, transient failures, and new observations trigger controlled rewrites, adding diagnostic TaskStars, creating fallbacks, rewiring dependencies, or pruning completed nodes, so the system adapts dynamically instead of aborting on errors wu2024agentkit.
  • Heterogeneous, asynchronous, and safe orchestration. Each TaskStar is matched to the most suitable device agent via rich AgentProfiles reflecting OS, hardware, and capabilities. The Constellation Orchestrator executes tasks asynchronously, allowing multiple TaskStars to progress in parallel. Safe assignment locking, event-driven scheduling, DAG consistency checks, and batched edits collectively ensure correctness and concurrency safety, achieving high efficiency without compromising reliability. These guarantees are further reinforced through formal verification.
  • Unified Agent Interaction Protocol (AIP). Built atop persistent WebSocket channels, we develop AIP, a protocol that provides a unified, secure, and fault-tolerant layer for agent registry, session management, task dispatch, and coordination. It ensures reliability under network fluctuations through automatic reconnection and retry, while exposing a lightweight, extensible interface that allows new agents to integrate seamlessly into the UFO3 ecosystem yang2025survey.
  • Template-driven framework for MCP-empowered device agents. To democratize agent creation, UFO3 provides a lightweight development template and toolkit for rapidly building new device agents. Developers can declare capabilities, bind to local environments, and extend them through one or more Model Context Protocol (MCP) servers hou2025model for tool augmentation. This modular design accelerates integration while maintaining consistency across the constellation.
Together, these designs enable the system to decompose, schedule, execute, and adapt distributed tasks efficiently while maintaining safety and consistency.
Building on these designs, we implemented the full UFO3 system as a comprehensive distributed system implementation with over 73K lines of Python code. The implementation integrates all major components, the centralized ConstellationAgent, the asynchronous Constellation Orchestrator, the AIP communication layer, and representative device agents for Windows, Linux and Android (mobile), each designed for containerized deployment and cross-environment compatibility. The system follows a modular, plugin-oriented architecture with type-safe interfaces, persistent telemetry, and built-in tracing, enabling reproducible, large-scale orchestration across heterogeneous devices. We further provide a futuristic WebUI for operator interaction and system visualization.
We evaluated UFO3 on NebulaBench, a benchmark of 55 cross-device tasks spanning 10 categories across 5 machines (a Windows 11 desktop, three Ubuntu CPU hosts, and one Ubuntu A100 GPU node). UFO3 achieves a Subtask Completion Rate (SCR) of 83.3% and a Task Success Rate (TSR) of 70.9%. It exposes substantial parallelism, with an average execution width of 1.72 (peaking at $\sim$3.5), and reduces end-to-end latency by 31% compared to a sequential baseline. Fault-injection experiments further demonstrate its robustness: UFO3 automatically retries and migrates under transient outages, gracefully degrades under partial failures, and recovers conservatively under global failures.
In essence, UFO3 dissolves device boundaries and transforms the digital estate into a single, adaptive collaborator brudy2018investigating; marks2020multi; zhang2018towards. It unifies distributed devices into a cohesive digital organism, one that executes user intents safely, asynchronously, and efficiently across heterogeneous environments. At its core, the Agent Interaction Protocol (AIP) serves as the connective tissue of this ecosystem, roviding a unified, fault-tolerant, and extensible communication substrate that allows new agents to join seamlessly and interoperate reliably. Over time, multiple constellations can interconnect through AIP, weaving together agents, devices, and capabilities into a self-organizing Digital Agent Galaxy. Through this design, UFO3 redefines cross-device automation, elevating it from a brittle engineering challenge to a unified orchestration paradigm, where multi-device workflows become naturally expressive and scalable across the landscape of ubiquitous computing.
2 Background
2.1 Digital Agents
Leveraging the power of LLMs, modern digital agents have emerged as powerful interfaces bridging human intent and the complex software ecosystems people interact with daily zhang2024large. These agents can parse natural language requests, interpret screenshots and system state zheng2025vem; wu2025gui; zhao2025learning, decompose complex goals into subtasks huang2024understanding, and generate scripts or commands to execute tasks using a variety of tools wang2024survey. Execution modalities include API calls, operating-system-level commands, GUI interactions via automation or accessibility interfaces, and code generation zhang2025api; qiao2023taskweaver; wang2024large; zhang2025swe.
Digital agents operate across a wide spectrum of platforms. Web-based agents can navigate browsers and search the Internet ning2025survey; zheng2024gpt, mobile agents are embedded in smartphone applications to automate mobile tasks zhang2025appagent; wang2024mobile, and desktop or laptop agents interact with local operating systems and graphical user interfaces zhang2025ufo; zhang2025ufo2. Across these platforms, agents observe system state, reason about tasks, and execute actions, enabling applications such as automated customer support flows, desktop productivity macros, and repetitive workflow automation, essentially acting as intelligent digital assistants.
Despite their versatility and growing adoption, existing digital agents are fundamentally limited by their confinement to a single device or environment. In practice, many modern workflows are no longer confined to a single endpoint brudy2018investigating: users frequently interact with a combination of desktops, mobile devices, cloud services, and specialized hardware, all as part of a single task xu2021cross; chen2020multi. Examples include running data analysis on a GPU cluster, collecting results on a personal laptop, and generating visual summaries on a tablet, or coordinating cross-platform deployments that touch both local workstations and cloud infrastructure. This trend makes the ability to operate seamlessly across devices increasingly urgent and commonplace.
Single-device frameworks face several inherent challenges in meeting this demand:
  • Limited device capabilities. Each agent is restricted by the hardware and software environment of its host, which limits the scope of tasks it can execute independently.
  • Fragmented personalization and context. User-specific preferences, personalization, and contextual knowledge captured on one device are difficult to transfer or leverage on another cemri2025multi, preventing a coherent multi-device experience.
  • Manual coordination overhead. Orchestrating multiple single-device agents to accomplish cross-device workflows typically requires extensive manual development and careful sequencing, which is time-consuming, error-prone, and hard to maintain.
Together, these limitations highlight the pressing need for a new generation of digital agents that can natively reason about, orchestrate, and adapt across heterogeneous devices.
2.2 Cross-Device Agent: A New Paradigm
To overcome the limitations of single-device frameworks, we introduce the concept of cross-device intelligence, in which digital agents collaborate seamlessly across multiple heterogeneous endpoints brudy2019cross. In this metaphorical digital galaxy, each device, desktop, mobile, cloud service, or specialized hardware, acts as a star, and coordinated tasks form a constellation that collectively fulfills complex user requests. Unlike traditional agents confined to a single host, these cross-device agents can reason about device capabilities, orchestrate subtasks, and execute actions as part of a unified, intelligent ecosystem cheninternet
Cross-device intelligence enables a qualitatively new user experience. A single natural-language request, such as “prepare a production-ready demo, run performance tests, and generate a report”, can trigger an orchestrated constellation of actions: builds executed on developer laptops, computation-heavy tests on GPU clusters, deployments on cloud containers, and visualization or report generation on local or mobile devices. To the user, this appears as one coherent, effortless operation, eliminating the need for manual coordination and device-specific intervention.
This paradigm directly addresses the key limitations of single-device agents. It overcomes the constraints of individual devices by assigning tasks to the endpoints with the necessary capabilities and resources. It allows personalization, context, and user preferences captured on one device to propagate across others, enabling a coherent multi-device experience giusti2025federation. And it automates orchestration of complex workflows, removing reliance on brittle glue code or labor-intensive manual integration.
2.3 Design Challenges in Cross-Device Orchestration
However, synchronous control loops, cross-device agent orchestration introduces a fundamentally different class of systems challenges. In a distributed, heterogeneous environment, agents must collaborate across device, network, and platform boundaries, each with distinct runtime contexts and execution semantics tran2025multi. Building a reliable and efficient orchestration layer under such conditions requires addressing three core challenges.
  1. 1.
    Asynchronous Parallelism. In cross-device workflows, multiple agents may execute concurrently on different endpoints. Unlike linear single-agent plans, task execution must accommodate partial completions, delayed feedback, and dynamic dependency resolution. The orchestrator must therefore reason about concurrency, detect when subtasks can safely proceed in parallel, and continuously adapt its scheduling plan based on evolving runtime states and network latencies yu2025dyntaskmas. Failing to do so may lead to wasted computation or stalled progress.
  2. 2.
    Distributed Coordination. Agents in a constellation operate across diverse network and trust domains, implemented in different languages and deployed on various infrastructures. Achieving coherent coordination among them requires standardized mechanisms for registration, capability advertisement, task dispatch, and result collection. These operations must occur through persistent, low-latency channels that can tolerate temporary disconnections and guarantee consistent task states cheninternet. In practice, ad-hoc HTTP calls or ephemeral connections are insufficient; a structured communication substrate is essential to sustain long-lived agent interactions.
  3. 3.
    Heterogeneous Extensibility. A cross-device ecosystem must embrace diversity rather than constrain it. Device agents differ in operating systems, execution environments, and available toolchains. The architecture should therefore allow rapid development, deployment, and integration of new device agents while ensuring consistent semantics in task execution and error propagation wu2024autogen. This requires a flexible interface model and a unified protocol to abstract platform-specific complexity, enabling scalable and evolvable orchestration across an open agent federation yang2025survey.
These challenges are not merely engineering details, as they delineate the boundary between ad-hoc agent scripts and a principled, distributed orchestration system. To address them, we present UFO3, a reliable and scalable framework for intelligent cross-device agent orchestration. UFO3 unifies four essential design elements: (i) a global ConstellationAgent that decomposes user intents into dynamic, dependency-aware task DAGs (Section 5); (ii) an event-driven orchestration engine that enables concurrent, asynchronous, and adaptive execution (Section 6); (iii) a standardized Agent Interaction Protocol (AIP) that supports persistent, low-latency, and extensible communication across heterogeneous agents (Section 7); and (iv) an easy-to-use development interface that allows developers to quickly build new device agents and seamlessly integrate them into the ecosystem (Section 8). Together, these mechanisms transform isolated device agents into a coherent, cooperative constellation capable of executing complex, distributed tasks efficiently, safely, and at scale.
3 UFO3: Design Principles and Overview
Figure 2: Layered architecture of UFO3.
We now present an overview of the UFO3 architecture, which transforms a collection of heterogeneous device agents into a unified, fault-tolerant execution fabric. Figure 2 depicts its layered design. At a high level, UFO3 follows a hierarchical orchestration model that separates global coordination from local execution. This separation enables scalable cross-device orchestration while maintaining consistent control and responsiveness across diverse operating systems and network environments.
3.1 Hierarchical Control Plane
At the top of the hierarchy, the ConstellationClient serves as the global control plane. It maintains a live registry of all connected device agents, including their capabilities, system specifications, and runtime health metrics. This registry allows the orchestrator to place tasks on devices that can satisfy their resource requirements, avoiding mismatches between task demands and device capacity.
Each device hosts a device agent server that manages local orchestration. The server maintains a persistent WebSocket session with the ConstellationClient and oversees execution contexts on the host. A lightweight device client on each host provides a unified interface to underlying tool environments, exposed via MCP servers, enabling task execution, telemetry streaming, and resource monitoring. This layered control plane cleanly decouples global orchestration policies from device-specific heterogeneity, providing a consistent abstraction across endpoints that may differ in OS, hardware, or network conditions.
3.2 Orchestration Flow
When a high-level user request arrives, the ConstellationClient invokes the ConstellationAgent to construct a TaskConstellation: a dynamic directed acyclic graph (DAG) that encodes task decomposition, dependencies, and candidate device mappings. Each node, or TaskStar, represents an atomic execution unit assigned to a suitable device agent according to its capability profile and current system load.
The Constellation Orchestrator executes the DAG asynchronously and in an event-driven manner. Task completions trigger dependent nodes, while failures prompt retry, migration, or partial DAG rewrites. This design allows workflows to adapt to real-time system dynamics, such as device churn, network variability, or incremental task updates, while preserving overall progress and global consistency. The result is an execution model that is both highly parallel and resilient, sustaining workflow completion even as subsets of devices fail or reconnect.
3.3 Cross-Agent Communication
All cross-agent interactions, including agent registration, capability synchronization, task dispatch, progress reporting, and result aggregation, are handled by the Agent Interaction Protocol (AIP). Built on persistent WebSocket channels, AIP provides a lightweight, bidirectional, and multiplexed substrate for structured event messages. Its design ensures low-latency propagation of control signals and consistent global state, even in the presence of intermittent connectivity or asynchronous updates.
Together, these design elements allow UFO3 to orchestrate large-scale, heterogeneous, and adaptive workflows, forming a cohesive foundation for building a resilient, multi-device execution fabric.
4 Formal Constellation Model
We now formalize the concept of a TaskConstellation, the central abstraction that captures the concurrent and asynchronous structure of distributed task execution and dependencies. This model provides the theoretical foundation for reasoning about task dependencies, execution order, and fault-tolerant orchestration across heterogeneous devices. At its core, a TaskConstellation represents a decomposed view of a complex user request in a directed acyclic graph (DAG) representation: a set of interdependent subtasks connected through explicit dependency edges. This formalism not only enables consistent scheduling and recovery but also supports runtime dynamism, allowing new tasks or dependencies to be introduced as the workflow evolves.
This representation provides clear advantages. Task ordering and dependencies are explicitly captured, ensuring correctness across distributed execution. The DAG topology naturally exposes parallelism and asynchronous execution, enabling efficient concurrency across heterogeneous devices. Moreover, nodes and edges can be dynamically added, removed, or rewired based on predecessor completion, allowing adaptive execution without compromising consistency. These properties make DAGs a natural and effective abstraction for modeling complex, cross-device workflows in heterogeneous environments.
4.1 TaskStar: Atomic Execution Unit
A TaskStar denotes the atomic unit of computation in the UFO3 framework, the smallest indivisible task scheduled on a device agent. Each TaskStar encapsulates the complete context necessary for autonomous execution, including its semantic description, assigned device, execution state, and dependency relationships.
  • Description: a natural-language specification of the task, sent to the target device agent;
  • Tips: A list of natural-language guidance designed to help the device agent successfully complete the task;
  • Device: the identifier of the device agent responsible for execution;
  • Status: the current execution state (e.g., pending, running, completed);
  • Dependencies: references to prerequisite tasks that must complete before execution.
Formally, a TaskStar $t_{i}$ is defined as:
$t_{i}=(\text{name}_{i},\text{description}_{i},\text{device}_{i},\text{tips}_{i},\text{status}_{i},\text{dependencies}_{i})$
Intuitively, a TaskStar “knows” what it should do, where it should run, how far it has progressed, and which other tasks it depends on. This self-contained representation enables fine-grained monitoring and decentralized scheduling across devices.
4.2 TaskStarLines: Dependency Edge
A TaskStarLine represents a dependency relation between two TaskStars, forming a directed edge in the task graph. Let $t_{i}$ and $t_{j}$ denote two TaskStars. A TaskStarLines $e_{i\rightarrow j}$ specifies that $t_{j}$ cannot begin until certain conditions on $t_{i}$ are satisfied:
$e_{i\rightarrow j}=(\text{from\_task}_{i},\text{to\_task}_{j},\text{type},\text{description})$
The type parameter determines the nature of the dependency:
  • Unconditional: $t_{j}$ always waits for $t_{i}$ to complete;
  • Success-only: $t_{j}$ proceeds only if $t_{i}$ succeeds;
  • Conditional: $t_{j}$ proceeds based on a user-defined or runtime condition.
These dependency edges enforce causal consistency within the constellation, ensuring that concurrent execution respects logical task ordering while maximizing parallelism.
4.3 TaskConstellation: Directed Acyclic Task Graph
A complete TaskConstellation is a DAG that encodes the structure of a distributed workflow:
$\mathcal{C}=(\mathcal{T},\mathcal{E})$
where $\mathcal{T}$ is the set of all TaskStars and $\mathcal{E}$ is the set of TaskStarLines. Each TaskConstellation provides a compact yet expressive representation of how a complex task decomposes into independently executable units with explicit dependencies.
Figure 3: Example of a TaskConstellation illustrating both sequential and parallel dependencies.
Figure 3 illustrates a simple TaskConstellation spanning multiple devices. Task A (LinuxAgent) must finish before Task C starts; both Task B (LinuxAgent) and Task C (WindowsAgent) must complete before Task D (WindowsAgent) begins; and Task E (MobileAgent) depends on the successful completion of both Task C and Task D. This example demonstrates how the model naturally captures both sequential and parallel dependencies within a unified structure.
4.4 Runtime Dynamism and Adaptivity
Unlike static DAG schedulers Islam2012Oozie; di2017nextflow; argo, UFO3 treats TaskConstellations as mutable objects. Tasks and dependency edges can be inserted, removed, or modified at runtime. This design enables UFO3 to dynamically react to evolving execution contexts, such as new user inputs, intermediate results, or device failures, without restarting the entire workflow. Such adaptivity is key to maintaining progress and efficiency in long-running, cross-device task orchestration.
In summary, the TaskConstellation formalism serves as the conceptual backbone of UFO3. It provides a precise and extensible representation of distributed workflows, enabling both rigorous reasoning and practical orchestration under asynchronous and failure-prone environments.
5 ConstellationAgent: The Centralized Constellation Weaver
Figure 4: An overview of the ConstellationAgent.
While the previous section formalized the TaskConstellation as an abstract model of distributed task structure, we now turn to its realization in the runtime control plane. The ConstellationAgent is the central intelligence of UFO3, responsible for interpreting user intent, constructing executable constellations, and steering their evolution across heterogeneous devices. Residing within the ConstellationClient, ConstellationAgent bridges the gap between high-level natural-language goals and concrete multi-agent execution through a unified orchestration interface.
As illustrated in Figure 4, ConstellationAgent acts as both a planner and a replanner: it first synthesizes a TaskConstellation from user instructions, then incrementally refines this constellation as feedback arrives from distributed agents. Internally, ConstellationAgent is implemented as an LLM-driven ReAct agent yao2022react governed by a finite-state machine (FSM) schneider2005state. This FSM alternates between two complementary operating modes, namely creation and editing, forming a closed-loop control cycle that continuously updates the global execution graph in response to runtime conditions.
This design achieves a tight coupling between symbolic reasoning and distributed execution: declarative goals are grounded into concrete task graphs, and dynamic feedback drives continuous graph mutation. Through this feedback-driven control loop, ConstellationAgent maintains global consistency, ensures forward progress, and adapts seamlessly to changing device conditions. Overall, the design provides three core benefits:
  1. 1.
    Unified reasoning and control: High-level task synthesis and low-level execution coordination are decoupled yet remain tightly synchronized via the TaskConstellation abstraction.
  2. 2.
    Dynamic adaptability: The editable TaskConstellation enables recovery, reallocation, and opportunistic task generation in the face of partial failures or evolving goals.
  3. 3.
    End-to-end observability: ConstellationAgent maintains a complete lineage of task states and dependencies, enabling introspection, debugging, and verifiable traceability.
5.1 Responsibilities and I/O Interface
ConstellationAgent serves as the reasoning core of the UFO3 control plane, orchestrating a structured feedback loop that alternates between creation and editing phases. Each phase defines explicit inputs, outputs, and operational responsibilities, ensuring that cross-device execution remains both consistent and adaptive.
Creation Mode.
In the creation phase, ConstellationAgent receives three primary inputs: (i) a user-issued goal, expressed in natural or structured language; (ii) the AgentProfile registry, describing each available device agent’s capabilities, environment, and metadata; and (iii) demonstration examples to support in-context learning (ICL) dong2022survey; jiang2024xpert.
Leveraging the LLM’s semantic reasoning capabilities, ConstellationAgent decomposes the user goal into a structured execution graph, the TaskConstellation. Each node (TaskStar) is annotated with explicit dependencies, resource constraints, and a target device assignment. This graph constitutes the initial execution plan, which is then handed off to the Constellation Orchestrator for distributed scheduling.
The TaskConstellation is generated in a structured, machine-readable JSON format. Beyond the raw graph, ConstellationAgent produces a detailed reasoning trace: it first formulates an Observation of the input and AgentProfile, followed by a structured Thought capturing its analysis and decision-making rationale wei2022chain; ding2024everything. It then outputs the next State in the control loop and the overall Result for user-request or the generated TaskConstellation. This ensures that the TaskConstellation is derived through a transparent, deliberate reasoning process.
Editing Mode.
During distributed execution, ConstellationAgent enters the editing phase of its adaptive control loop. In this mode, it continuously consumes: (i) the original user request; (ii) the current AgentProfile registry; (iii) a serialized snapshot of the TaskConstellation; and (iv) demonstration examples for ICL.
Upon receiving completion or failure events, ConstellationAgent evaluates whether modifications are necessary, such as adding follow-up subtasks, removing redundant ones, or refining dependency edges. Edits are applied only to non-terminal TaskStars (i.e., not in Running, Completed, or Failed), ensuring runtime correctness while allowing the TaskConstellation to evolve dynamically.
Editing actions are invoked through one or more of the tools defined in Section 5.3 to facilitate parsing and execution. Concurrently, ConstellationAgent produces a structured reasoning trace, including a Thought narrative that explains the current constellation state, justifies any modifications (or lack thereof), identifies the next State in the control loop, and summarizes the overall user-request Result. This transparent, step-by-step reasoning guarantees that adaptations are both explainable and consistent.
This dual-mode control pattern realizes a balance between global consistency and local adaptivity. By structuring feedback integration through explicit modes and invariants, ConstellationAgent achieves stable yet flexible orchestration, avoiding common pitfalls of uncontrolled LLM self-modification. The resulting system embodies a feedback-driven orchestration loop that remains both reactive and verifiable.
5.2 Finite-State Machine and Lifecycle
Figure 5: Lifecycle state transitions of the ConstellationAgent.
The internal logic of ConstellationAgent is expressed as a finite-state machine (FSM), providing a clear, enforceable structure for task lifecycle management. We show its state transition in Figure 5. The FSM governs how ConstellationAgent transitions across four primary operational states:
  • START: Initialization state where ConstellationAgent receives the user goal and constructs the initial TaskConstellation (creation mode).
  • CONTINUE: The steady-state loop, where task progress is monitored and incremental edits are applied based on runtime feedback (editing mode).
  • FINISH: The successful termination state, triggered when all subtasks are completed or no further edits are required. Results are aggregated and reported to the user.
  • FAIL: The terminal error state, entered upon irrecoverable failures or unreachable goals, prompting abort and logging for user recovery.
This FSM-based design ensures deterministic task transitions, consistent global state evolution, and predictable recovery paths. It also provides a natural boundary between LLM reasoning and deterministic control logic, which improves safety and debuggability in complex, cross-device workflows.
5.3 Constellation MCP Server: Structured Task Management
Table 1: Core tools exposed by the Constellation MCP Server for managing tasks and dependencies.
Tool Name Purpose Input Output
add_task Add a new atomic task (TaskStar) Task ID, name, description, target device, tips Updated TaskConstellation
remove_task Remove a task and all associated dependencies Task ID Updated TaskConstellation
update_task Modify task fields (name, description, device, tips) Task ID + updated fields Updated TaskConstellation
add_dependency Establish a dependency between two tasks Dependency ID, from task, to task, condition Updated TaskConstellation
remove_dependency Remove a dependency line Dependency ID Updated TaskConstellation
update_dependency Update the condition or description of a dependency Dependency ID, condition description Updated TaskConstellation
build_constellation Batch-create tasks and dependencies from structured input Configuration dictionary, clear flag Built TaskConstellation
To operationalize dynamic graph construction, ConstellationAgent interacts with a lightweight Constellation MCP Server hou2025model, a modular component exposing a standardized set of task and dependency management primitives. This server serves as the structured manipulation layer that bridges LLM-level reasoning and concrete execution state. Each operation encapsulates a single, idempotent transformation on the TaskConstellation, ensuring reproducibility and easy rollback.
The MCP Server supports both fine-grained (single task or edge) and bulk (batch graph) operations, all of which return a serialized, globally consistent constellation snapshot. Table 1 summarizes the core toolset. Through this uniform interface, ConstellationAgent can safely evolve task graphs at runtime while preserving key invariants such as DAG validity and single assignment.
By decoupling reasoning (handled by ConstellationAgent) from structured mutation (handled by MCP), UFO3 achieves a clean separation of concerns: ConstellationAgent focuses on semantic decision-making, while the MCP Server enforces syntactic and structural integrity. This design not only enhances robustness and auditability but also facilitates future extensibility, for instance, incorporating new agent types, validation hooks, or task optimization heuristics without modifying the orchestration core.
Summary.
In summary, ConstellationAgent serves as the “central weaver” of distributed intelligence within UFO3. By combining LLM-driven reasoning, a finite-state control backbone, and a structured task manipulation interface, it transforms abstract user goals into live, evolving constellations, maintaining both rigor and adaptability across the lifecycle of multi-device orchestration.
6 Asynchronous Dynamic Constellation Orchestrator
Figure 6: The Constellation Orchestrator bridges TaskConstellation and execution, enabling asynchronous, adaptive task orchestration across devices.
With the ConstellationAgent governs the reasoning and evolution of a TaskConstellation, the Constellation Orchestrator brings that plan to life by executing, monitoring, and adapting interdependent tasks across heterogeneous devices li2022dag, as shown in Figure 6. Conceptually, it transforms a static DAG into a living execution fabric, where tasks evolve concurrently, react to runtime signals, and adapt to new decisions generated by the reasoning agent.
Unlike traditional serial or static agent workflows, UFO3’s orchestration must satisfy three often conflicting goals: (i) asynchronous parallelism to leverage device heterogeneity, (ii) safety and consistency under concurrent DAG updates, and (iii) adaptivity to runtime feedback from both devices and LLM reasoning. Achieving these goals poses several technical challenges. First, subtasks must be assigned and executed asynchronously without violating data dependencies. Second, task execution may overlap with live TaskConstellation edits, requiring strict consistency control. Third, edited graphs must remain valid and acyclic. Finally, frequent updates should not degrade performance or introduce synchronization bottlenecks.
To address these challenges, the Constellation Orchestrator is built around five design pillars: (1) event-driven coordination, (2) asynchronous scheduling, (3) safe assignment locking, (4) consistency enforcement, and (5) batched constellation editing. Together, these principles enable scalable, adaptive orchestration over evolving task graphs while preserving correctness and efficiency.
6.1 Event-Driven Coordination
Traditional DAG schedulers rely on polling or global checkpoints to detect task completion, introducing latency and synchronization overhead. In contrast, the Constellation Orchestrator operates as a fully event-driven system built on an internal event bus and an observer design pattern.
Two key components manage this process: the ConstellationProgressObserver, which tracks task execution and orchestrates DAG edits, and the ConstellationModificationSynchronizer, which ensures global consistency after each modification. The orchestrator emits and reacts to four primary event types:
  1. 1.
    TASK_STARTED: triggered when a TaskStar is assigned to a device and begins execution.
  2. 2.
    TASK_COMPLETED: emitted upon successful task completion, prompting potential DAG updates.
  3. 3.
    TASK_FAILED: emitted on failure, triggering re-planning or fallback logic.
  4. 4.
    CONSTELLATION_MODIFIED: emitted once DAG edits are committed and synchronized across agents.
These events collectively capture the lifecycle of each task and the evolution of the constellation. All handlers operate asynchronously, ensuring immediate, fine-grained reactions to runtime signals without centralized coordination delays. This event-driven design provides high responsiveness and forms the foundation of adaptive orchestration in UFO3.
6.2 Asynchronous Scheduling
Figure 7: Illustration of asynchronous scheduling and concurrent TaskConstellation editing.
At the core of the orchestrator lies a fully asynchronous scheduling loop. Unlike traditional schedulers that alternate between discrete planning and execution phases, the orchestrator continuously monitors the evolving DAG to identify ready TaskStar, those whose dependencies are satisfied, and dispatches them concurrently to available devices. Each TaskStar runs within an asyncio coroutine that encapsulates its full lifecycle: execution, result collection, and event publication. When a task starts, a TASK_STARTED event is emitted; upon completion or failure, corresponding TASK_COMPLETED or TASK_FAILED events are immediately published to trigger downstream orchestration updates.
Notably, TaskStar execution and TaskConstellation editing can proceed concurrently. As shown in Figure 7, when Task A completes and triggers an edit, the edit operation executes in parallel with the ongoing Tasks B and C. Similarly, edits triggered by Task B overlap with subsequent executions. This concurrency further reduces end-to-end latency by overlapping computation and orchestration, allowing the system to adapt in real time as results stream in.
This asynchronous design maximizes device utilization and eliminates idle waiting, enabling parallel progress across heterogeneous devices. It is essential for scaling cross-device workflows, where independent subtasks (e.g., log collection, file aggregation, or model execution) can execute concurrently while higher-level orchestration continuously adapts to dynamic task states and outcomes.
6.3 Safe Assignment Locking and Synchronization
Input: Event stream $\mathcal{E}$, current TaskConstellation $\mathcal{C}$
Output: Consistent and updated $\mathcal{C}$ with newly scheduled ready tasks
1 while system is running do
2 foreach event $e\in\mathcal{E}$ do
3    if $e$ is TASK_COMPLETED or TASK_FAILED then
       async enqueue($e$) ;
       //record completion/failure for processing asynchronously
4
5
6
  acquire(assign_lock) ;
 //suspend new assignments
7
8 while queue not empty do
    $e\leftarrow$ dequeue() ;
    //get next event for processing
9
    $\Delta$ = invoke(ConstellationAgent, edit($\mathcal{C}$, $e$)) ;
    //propose DAG edits
    $\mathcal{C}\leftarrow$ apply($\mathcal{C}$, $\Delta$) ;
    //update the constellation structure
      validate($\mathcal{C}$) ;
    //ensure acyclicity and invariants (I1–I3)
      publish(CONSTELLATION_MODIFIED, $t$) ;
    //notify DAG update
    $\mathcal{C}\leftarrow$ synchronize($\mathcal{C}$, $\mathcal{T}_{C}$) ;
    //merge newly completed TaskStars
10
11
  release(assign_lock) ;
 //resume orchestration after all queued events are processed
12
13 // Rescheduling Phase (outside lock)
 $\mathcal{T}_{R}\leftarrow$ get_ready_tasks($\mathcal{C}$) ;
 //collect newly ready TaskStars
14 foreach $t\in\mathcal{T}_{R}$ do
    async dispatch($t$) ;
    //send to available device agent asynchronously
    async publish(TASK_STARTED, $t$) ;
    //notify asynchronously
15
16
Algorithm 1 Safe Assignment Locking and Asynchronous Rescheduling Protocol
Figure 8: An example of the safe assignment locking and event synchronization workflow.
While asynchrony improves efficiency, it introduces correctness challenges when task execution overlaps with DAG updates. Specifically, the orchestrator must prevent race conditions when the ConstellationAgent dynamically adds, removes, or rewires TaskStar during execution. Without safeguards, a task could be dispatched based on a stale DAG, leading to duplicated or invalid execution.
To ensure atomicity, the orchestrator employs a safe assignment lock. When an edit cycle begins, bounded by a TASK_COMPLETED/TASK_FAILED event and its corresponding CONSTELLATION_MODIFIED event, the scheduler suspends new task assignments to prevent dispatching based on stale DAG states. During this period, incoming events are queued, and the ConstellationModificationSynchronizer guarantees that all edits are applied atomically. A key aspect of this process is synchronization: once the ConstellationAgent finishes editing, the orchestrator merges its structural changes with runtime updates from concurrently running tasks (i.e., new completions or failures that occurred during the editing window). This ensures that the final TaskConstellation reflects a globally consistent view of both reasoning-time modifications and execution-time progress.
[…]
This mechanism guarantees that task assignments remain immutable during edits, preventing conflicts between execution and modification. By maintaining atomicity and synchronization without blocking overall progress, it preserves both safety and consistency for concurrent, LLM-driven orchestration at scale.
6.4 Consistency and Safety Guarantees
Since the DAG may be dynamically rewritten by an LLM, the orchestrator enforces runtime invariants to preserve correctness even under partial or invalid updates:
  • I1 (Single Assignment): Each TaskStar has at most one active device assignment at any time.
  • I2 (Acyclic Consistency): Edits must preserve DAG acyclicity; the orchestrator performs local cycle detection before committing modifications.
  • I3 (Valid Update): Only PENDING tasks and their dependent nodes may be modified; RUNNING, COMPLETED, and FAILED nodes are immutable.
Together, these invariants ensure that even as the constellation evolves, new stars form, old ones fade, the overall structure remains stable and semantically valid. We enforce three runtime invariants (I1–I3) under a lock-bounded editing regime; see Appendix A for the formal state model and proof sketch.
6.5 Batched Constellation Editing
Frequent LLM-driven edits can introduce significant overhead if processed individually. To balance responsiveness with efficiency, the orchestrator supports batched constellation editing. During a reasoning round, multiple TASK_COMPLETED or TASK_FAILED events may accumulate; instead of invoking the ConstellationAgent after each event, the orchestrator aggregates them and applies the resulting modifications atomically once reasoning is complete. As illustrated in Figure 8, when task_A completes ($t_{0}$), the ConstellationAgent starts an edit cycle. During this process, new completion events from task_B and task_C arrive ($t_{3}$, $t_{4}$) and are temporarily queued. After the first edit result $\Delta_{A}$ is validated and synchronized ($t_{5}$), the orchestrator batches the pending updates from B and C into a single reasoning round ($t_{6}$$t_{7}$).
This batching mechanism amortizes LLM invocation and synchronization overhead while preserving atomicity and consistency. It enables the orchestrator to remain both efficient and adaptive, reacting swiftly to meaningful state transitions without incurring excessive micro-edits or redundant reasoning calls. We prove an edit–sync confluence lemma showing that folding runtime events commutes with lock-bounded edits within the same window; see Appendix A (Edit–Sync Confluence).
Design Summary.
In contrast to static DAG or synchronous schedulers, the Constellation Orchestrator treats task execution as an open-world process, continuously evolving, reacting, and converging toward user intent. Its event-driven backbone ensures responsiveness; asynchronous scheduling maximizes concurrency; locking and batching ensure safety; and DAG validity checks preserve correctness under dynamic reasoning.
Together, these components realize a new form of orchestration, asynchronous, adaptive, and reasoning-aware, that bridges declarative intent and distributed execution across a heterogeneous universe of intelligent agents.
7 Agent Interaction Protocol (AIP)
The orchestration model described in Section 6 requires a communication substrate that remains correct under continuous DAG evolution, dynamic agent participation, and fine-grained event propagation. Legacy HTTP-based coordination approaches (e.g., A2A duan2025agent, ACP ehtesham2025survey) assume short-lived, stateless interactions, incurring handshake overhead, stale capability views, and fragile recovery when partial failures occur mid-task. These assumptions make them unsuitable for the continuously evolving workflows and long-running reasoning loops characteristic of UFO3.
7.1 Design Overview
AIP serves as the nervous system of UFO3, connecting the ConstellationClient, device agent services, and device clients under a unified, event-driven control plane, as shown in Figure 9. It is designed as a lightweight yet evolution-tolerant protocol to satisfy six goals:
  1. (G1)
    Maintain persistent bidirectional sessions to eliminate per-request overhead;
  2. (G2)
    Unify heterogeneous capability discovery via multi-source profiling;
  3. (G3)
    Ensure fine-grained reliability through heartbeats and timeout managers for disconnection and failure detection;
  4. (G4)
    Preserve deterministic command ordering within sessions;
  5. (G5)
    Support composable extensibility for new message types and resilience strategies;
  6. (G6)
    Provide transparent reconnection and task continuity under transient failures.
To meet these requirements, AIP adopts a persistent, bidirectional WebSocket transport and decomposes the orchestration substrate into five logical strata (Figure 9), each responsible for a distinct aspect of reliability and adaptability:
  • L1: Message Schema Layer — Defines strongly-typed, Pydantic-validated contracts (ClientMessage, ServerMessage) for message direction, purpose, and task transitions. Structured metadata (system info, capabilities) supports unified capability discovery (G2) and deterministic ordering via explicit ID correlation (G4).
  • L2: Transport Abstraction Layer — Provides a protocol-agnostic Transport interface with a production-grade WebSocket implementation supporting configurable pings, timeouts, and large payloads. Decoupled transport logic ensures low-latency persistent sessions (G1) and future extensibility (G5).
  • L3: Protocol Orchestration Layer — Implements modular handlers for registration, task execution, heartbeat, and command dispatch (see Appendix B), each extending a common AIPProtocol base with middleware hooks (logging, metrics, auth). This design ensures ordered state transitions (G4) and composable extensibility (G5).
  • L4: Resilience and Health Management Layer — Encapsulates HeartbeatManager, TimeoutManager, and ReconnectionStrategy with exponential backoff and automatic session recovery. It guarantees reliability (G3) and seamless task continuity under transient disconnections (G6).
  • L5: Endpoint Orchestration Layer — Provides role-specific facades: DeviceServerEndpoint, DeviceClientEndpoint, and ConstellationEndpoint, integrating lower layers into deployable components. These endpoints unify connection lifecycle, task routing, and health monitoring across roles, reinforcing G1G6.
Together, these layers form a vertically integrated stack where L1 establishes semantic contracts, L2 provides transport flexibility, L3 implements protocol logic, L4 ensures operational resilience, and L5 delivers deployment-ready orchestration primitives. This design enables UFO3 to maintain correctness and availability under DAG evolution (G4, G5), agent churn (G3, G6), and heterogeneous execution environments (G1, G2).
L1: Message Schema LayerClientMessage, ServerMessage, Pydantic Validation L2: Transport Abstraction LayerWebSocketTransport, Persistent Connections, Adapters L3: Protocol Orchestration LayerRegistration, TaskExecution, Heartbeat, DeviceInfo, Command L4: Resilience & Health Management LayerHeartbeatManager, TimeoutManager, ReconnectionStrategy L5: Endpoint Orchestration LayerDeviceServer, DeviceClient, Constellation Endpoints ConstellationClient Device Agent Server Device Agent Client
Figure 9: AIP Architecture: Five-layer protocol stack enabling persistent, resilient multi-agent orchestration.
7.2 Agent Registration and Profiling
Agent registration in AIP corresponds to the entry point of the orchestration pipeline, anchoring the capability discovery and topology formation processes outlined in (G2) and implemented primarily through the L1–L3 layers.
Figure 10: Agent registration flow: multi-source AgentProfile construction and registration.
As illustrated in Figure 10, the registration pipeline consists of three complementary stages that together establish a unified and continuously refreshed view of the constellation’s capabilities:
  1. 1.
    User-specified registration (ConstellationClient). Administrators provide endpoint identities and specify user preferences. These initial configurations define the logical boundaries of the constellation and seed the connection parameters required for session establishment.
  2. 2.
    Service-level manifest (device agent service). Each device agent advertises its supported tools, environment variables, and operational metadata through a REGISTER message. These descriptors are validated and normalized by the Message Schema Layer (L1) and merged through the Protocol Orchestration Layer (L3), ensuring semantic consistency and structured capability discovery.
  3. 3.
    Client-side telemetry (device agent client). Local clients continuously report runtime metrics, such as OS version, hardware status, GPU utilization, and software environment, to the device agent service. This telemetry stream keeps the global registry up to date and enables adaptive re-scheduling under resource drift or device churn.
AgentProfile: gpu_agentStatus: idle  System
OS: linux  (Ubuntu 24.04)
Capabilities: cli, file_system, data_processing, model_training
Paths
Dataset: ~/dataset
Training: ~/script
Code: ~/code
Performance
GPU: 4 $\times$ NVIDIA A100 80G
CPU: 96 cores
Memory: 866.1 GB
Network & Host (optional)
Host: gui-model-a100
IP: 172.19.0.4
  Last heartbeat: 2025-10-28 07:46:24 UTC  |   Server: ws://localhost:5005/ws
Figure 11: An example AgentProfile of a GPU agent with Linux system.
The aggregated results from these three stages are merged by the ConstellationClient into a unified AgentProfile that represents each agent’s real-time operational state. An example AgentProfile for a GPU-enabled device is shown in Fig. 11, demonstrating how multi-level profiling captures hardware, software, and dynamic runtime descriptors within a single schema.
Through this registration and profiling pipeline, AIP achieves continuous capability discovery, evolution-tolerant orchestration, and consistent topology awareness across heterogeneous devices. The process directly fulfills (G1–G3) by maintaining persistent sessions, ensuring reliable metadata propagation, and enabling transparent adaptation as the constellation evolves.
7.3 Task Dispatch and Result Delivery
Task dispatch operationalizes the event-driven execution model envisioned in (G1) and (G4) through tightly managed sessions that persist across multiple task rounds. When the ConstellationClient assigns a TaskStar to a device, the Transport Abstraction Layer guarantees low-latency delivery of the serialized TASK message to the target agent service. The Protocol Orchestration Layer coordinates message routing, while the Resilience and Health Management Layer monitors the session heartbeat to ensure reliability.
Each task follows a deterministic life cycle, from TASK to TASK_END, with strict ordering guarantees enforced by the session-level sequence manager. Intermediate logs and evaluator outputs are streamed back incrementally to the ConstellationClient, which updates the global TaskConstellation state and triggers potential DAG adjustments.
This continuous, feedback-driven execution loop transforms AIP from a mere transport protocol into a temporal coordination substrate, harmonizing asynchronous reasoning, scheduling, and execution across distributed devices.
7.4 Command Execution
At a finer operational granularity, AIP implements a unified command execution model that directly fulfills (G4) and (G5) by ensuring deterministic, extensible control within persistent sessions.
Each COMMAND message specifies a unique identifier, target function, and a typed argument list. The Message Schema Layerenforces structure and validation, while the Protocol Orchestration Layer executes commands sequentially within the session context to maintain determinism. To optimize multi-action workflows, multiple commands may be batched in a single message, reducing round-trip overhead. Execution results are returned as structured Result objects, containing status codes, return values, and error metadata. Failures or timeouts are propagated through the same channel, enabling the orchestrator to apply adaptive recovery or task reassignment strategies via L4’s resilience mechanisms.
Unified with UFO3’s MCP tool-calling interface, this model bridges system-level orchestration with model-generated actions, ensuring that high-level reasoning and low-level execution operate under a consistent and evolvable protocol surface.
7.5 Resilient Connection Protocol
The distributed and volatile nature of device environments necessitates a dedicated resilience layer to uphold (G3) and (G6). AIP’s Resilient Connection Protocol, implemented primarily within L4, guarantees synchronized fault handling and seamless recovery across client–server boundaries.
When a Device Agent disconnects unexpectedly, the orchestrator immediately marks it as DISCONNECTED, removes it from the active scheduling pool, and triggers background reconnection attempts using exponential backoff. Upon recovery, the agent re-registers automatically, restoring its prior session state and resuming task participation without manual intervention. If disconnection occurs during task execution, all affected tasks are transitioned to TASK_FAILED, and corresponding updates are propagated to the ConstellationAgent for DAG revision, ensuring that the orchestration view remains globally consistent.
Symmetrically, when the ConstellationClient itself disconnects, the corresponding Device Agent Server receives a termination signal and proactively aborts all ongoing tasks associated with that client. This bidirectional fault-handling policy prevents resource leakage, avoids orphaned execution states, and guarantees that both endpoints maintain a consistent global view of task progress.
Together, these mechanisms realize an end-to-end resilient orchestration substrate that preserves correctness, availability, and synchronization under transient network failures or partial system outages,closing the reliability loop envisioned in AIP’s layered design.
Summary.
AIP consolidates registration, task dispatch, command execution, and resilience into a coherent, evolution-tolerant communication fabric that embodies the six goals outlined in Section 7.1. Functionally, AIP forms the nervous system of UFO3, enabling reasoning, execution, and recovery to operate seamlessly within an evolving constellation of intelligent agents. Its minimal yet extensible design ensures that as workflows, models, and environments evolve, the underlying communication protocol remains stable, adaptive, and correct by construction.
8 Design and Development of Device Agents
With the structured reasoning of the ConstellationAgent, the dynamic execution model of the Constellation Orchestrator, and the low-latency, persistent communication enabled by AIP, the next challenge is clear: how do we design a device agent that can be quickly onboarded into UFO3, adapt to a heterogeneous platform, and seamlessly participate in evolving task constellations? Our solution provides a standardized template for device agent development, minimizing engineering effort while maximizing system scalability and reliability.
Figure 12: The three-layer framework of a device agent.
8.1 Architecture Overview
A device agent serves as the execution endpoint of UFO3, translating high-level task directives into concrete actions on a target device. To support rapid development and flexible deployment, each agent is structured as a three-layer, state-machine-driven framework, illustrated in Figure 12. To enable safe and scalable execution across heterogeneous devices, each agent is further partitioned into a server that manages orchestration and FSM logic, and a client that executes low-level commands locally. This server-client separation, combined with the layered FSM design, balances modularity, extensibility, and runtime robustness, and supports both single-agent and multi-agent deployment scenarios depending on platform requirements.
Specifically, the architecture decomposes agent behavior into three hierarchical levels:
  1. 1.
    Level-1: State (Finite-State Machine Layer). Each agent maintains an internal state machine that governs its behavior at each execution step. A state encapsulates a processor, the next state to transition to, and optionally the next agent to invoke in multi-agent setups. The collection of states defines a finite-state machine that ensures predictable, controllable lifecycle progression.
  2. 2.
    Level-2: Strategy (Execution Logic Layer). Within each state, the processor manages a sequence of strategies that implement the step-level workflow. Strategies handle tasks such as data collection, environment inspection, prompt construction, action planning, or tool invocation. This separation allows the agent to compose complex behaviors from modular, reusable strategies, while maintaining clear boundaries between decision logic and execution mechanics.
  3. 3.
    Level-3: Command (System Interface Layer). Each strategy can invoke a set of commands from a configured MCP server, which provides standardized operations for perceiving system state, executing tools, or interacting with device resources. Commands are executed deterministically and report structured outcomes, allowing higher layers to react and adapt without managing low-level device specifics.
By instantiating an agent with concrete State, Strategy, and Command definitions, a fully functional device agent can be realized. This hierarchical, layered approach offers several advantages. First, the same framework can accommodate a wide variety of devices, platforms, and execution environments. Second, developers can onboard new devices by defining only the relevant states, strategies, and commands, without rewriting orchestration or communication logic. Finally, the layered FSM structure allows the agent to respond to dynamic task edits, partial failures, or concurrent executions while maintaining correctness.
Together, this architecture positions device agents as plug-and-play execution units, seamlessly bridging the high-level reasoning of the ConstellationAgent, the dynamic orchestration of the Constellation Orchestrator, and the event-driven communication of AIP. It forms the final, essential layer that enables UFO3 to operate as a cohesive, scalable, and resilient multi-device system.
8.2 Server-Client Architecture
Figure 13: The server-client architecture of a device agent.
To support safe, scalable, and flexible execution across heterogeneous devices, each device agent is partitioned into a server and a client, as illustrated in Figure 13. This separation of responsibilities aligns naturally with the layered FSM architecture and leverages AIP for reliable, low-latency communication.
Server: Orchestration and State Management.
The agent server is responsible for managing the agent’s state machine lifecycle, executing high-level strategies, and interacting with the ConstellationAgent or the orchestrator. It handles task decomposition, prompt construction, decision-making, and command sequencing. Crucially, the server maintains full control over the agent’s workflow logic, enabling updates to decision strategies without impacting low-level execution on the device.
Each server instance exposes its AgentProfile, a structured description of its capabilities, configurations, and runtime status. This metadata allows the orchestrator to dynamically select suitable agents for specific subtasks, improving task distribution efficiency. A single server can manage multiple agent clients concurrently, maintaining isolation across devices while supporting centralized supervision and coordination.
Client: Commands Execution and Resource Access.
The agent client runs on the target device and manages a collection of MCP servers or tool interfaces. These MCP servers can operate locally (via direct invocation) or remotely (through HTTP requests), and each client may register multiple MCP servers to access diverse tool sources. Upon receiving commands from the agent server, such as collecting telemetry, invoking system utilities, or interacting with hardware components, the client translates them into MCP tool calls, executes them deterministically, aggregates the results, and returns structured outputs via AIP. The client remains stateless with respect to reasoning: it faithfully executes directives without engaging in high-level decision-making.
During initialization, each client connects to the agent server through the AIP endpoint, performs self-checks (e.g., disk, CPU, memory, GPU, and network configuration), and registers its hardware–software profile. This profile is integrated into the server’s AgentProfile, giving the orchestrator complete visibility into system topology and resource availability for informed task assignment and scheduling.
Server-Client Communication.
All communication between the server and client is routed through the AIP, leveraging persistent WebSocket connections. This allows bidirectional, low-latency messaging that supports both synchronous command execution and asynchronous event reporting. By decoupling high-level reasoning from low-level execution, the system can safely update server logic or client MCP tools independently, without disrupting ongoing workflows.
Design Consideration.
This server–client architecture provides strong modularity and scalability. Device clients can be rapidly deployed with minimal setup, immediately joining UFO3 as execution endpoints. The server focuses on high-level reasoning and orchestration, while clients ensure deterministic command execution, preventing cross-layer interference and simplifying maintenance. Persistent sessions and structured AIP event semantics enhance robustness under intermittent connectivity and dynamic task updates. The design also scales efficiently across devices: a single server can orchestrate multiple clients, and extensibility is achieved by adding new tools or interfaces at the client side or new reasoning strategies at the server without mutual dependencies.
8.3 State: Finite-State Machine Layer
1class AgentState(ABC):
2 """Abstract interface for a device agent state."""
3
4 @abstractmethod
5 async def handle(self, agent, context=None):
6 """Execute the logic for the current state."""
7 pass
8
9 @abstractmethod
10 def next_state(self, agent) -> "AgentState":
11 """Return the next state in the FSM."""
12 pass
13
14 @abstractmethod
15 def is_round_end(self) -> bool:
16 """Determine whether the current task round has ended."""
17 pass
Figure 14: Simplified interface of a device agent state, highlighting the core methods for execution, state transition, and termination checks.
The top-level lifecycle of each device agent is governed by a finite state machine (FSM), which provides a structured and predictable execution framework. Each state encapsulates the logic for handling a specific step of the agent’s workflow, including invoking the corresponding processor, determining the next state to transition to, selecting the next agent in multi-agent setups, and deciding whether the current task has reached completion. Figure 14 illustrates the interface exposed by a state.
At runtime, the agent invokes the state’s handle function, which executes the state-specific behavior (e.g., a processor) and returns control decisions to the FSM. Transitions between states can be determined dynamically by the agent based on LLM reasoning, or triggered by rule-based logic in response to errors, timeouts, or external events. This flexibility allows the agent to react promptly to runtime conditions, while maintaining a predictable execution path for normal workflows.
The FSM-based design offers several advantages. First, it enforces a clear separation of concerns: state transitions govern workflow progression, processors implement step-level strategies, and commands handle low-level system interactions. Second, it simplifies reasoning about agent behavior and facilitates debugging, since each state represents an isolated, testable unit. Finally, the FSM enables device agent to safely manage dynamic agent behaviors and failures, and concurrent executions.
In essence, the State layer provides a robust backbone for device agent execution, allowing high-level orchestration from the ConstellationAgent and Constellation Orchestrator to be reliably translated into stepwise, adaptive actions across heterogeneous devices.
8.4 Strategy: Composable Execution Logic
Each agent state delegates step-level workflow management to a processor, which orchestrates a sequence of strategies. Strategies encapsulate modular execution logic, enabling fine-grained control over the agent’s behavior while maintaining a clear separation between decision-making (State layer) and concrete actions (Command layer).
In a typical processor, we define four core strategy types and execute sequentially:
  • DATA_COLLECTION: Gather necessary context from the device, such as screenshots, accessibility information, system status, or user input.
  • LLM_INTERACTION: Construct prompts using the collected data and interact with the LLM to obtain actionable instructions or decisions.
  • ACTION_EXECUTION: Perform the commands returned by the LLM or pre-defined toolkits, applying them to the device environment deterministically.
  • MEMORY_UPDATE: Update the agent’s short-term or long-term memory to reflect task progress and provide context for subsequent steps.
This strategy-based design offers several advantages. First, it allows flexible customization for different device types and task requirements; strategies can be reordered, added, or replaced without modifying the core state machine. Second, it provide a template that modularizes execution, making the workflow easier to test, debug, and extend. Finally, by clearly delineating data collection, reasoning, action, and memory update, the processor ensures that each step in the agent’s lifecycle is composable, observable, and adaptable to dynamic runtime conditions.
Overall, the Strategy layer acts as the operational bridge between the high-level reasoning of the State layer and the low-level Command execution, enabling device agents to carry out complex, multi-step tasks reliably and efficiently.
8.5 Command: Atomic Execution Units
Commands represent the atomic execution units within a device agent. Each command encapsulates a specific operation, defined by a function and its corresponding arguments, which maps directly to a tool call on the MCP server co-located with the device client. By treating commands as self-contained units, the agent can systematically decompose complex workflows into discrete, testable actions.
Each Strategy invokes commands to realize its operational intent, for example, a DATA_COLLECTION strategy might request a screenshot or accessibility tree, while an ACTION_EXECUTION strategy triggers a system command or UI interaction. Commands are transmitted via the AIP to the client, which performs the actual tool execution on the device and returns structured results back to the server. This separation of command logic from device-level execution ensures deterministic, reliable, and auditable operations across heterogeneous platforms. At runtime, the agent can query the client for available commands and their usage metadata, enabling dynamic selection by the LLM and adaptive workflows. This design supports extensibility: new tools or device capabilities can be integrated by simply registering additional commands at the client layer, without modifying the server-side logic or State/Strategy definitions.
In essence, the Command layer completes the device agent’s execution pipeline: it bridges high-level reasoning (State), step-wise workflow orchestration (Strategy), and concrete device interaction, providing a robust and flexible foundation for reliable, multi-step automation across diverse environments.
8.6 Example Device Agents: LinuxAgent and WindowsAgent
To demonstrate how the layered device agent architecture and server-client design can be instantiated in practice, we present two representative case studies: LinuxAgent and WindowsAgent. These examples illustrate how UFO3’s templates enable rapid development, integration, and execution of device agents across different platforms.
The LinuxAgent showcases a single-agent deployment, highlighting how a standalone agent can leverage the FSM-based State, Strategy, and Command layers to interact with system tools, collect telemetry, and execute workflow tasks. In contrast, the WindowsAgent demonstrates a multi-agent deployment, where multiple agent instances coordinate via a server-client setup to manage complex workflows involving local MCP servers, UI automation, and dynamic LLM-driven task decomposition. Together, these two case studies provide concrete examples of how UFO3 can accommodate diverse execution environments, and serve as the foundational agents for the experimental evaluations in Section 10.
8.6.1 LinuxAgent: Single-Agent CLI Execution
The LinuxAgent is designed as a lightweight, single-agent instance capable of executing command-line instructions to fulfill user requests. It demonstrates how a standalone device agent can leverage the layered FSM architecture and server-client design to perform intelligent, iterative task execution on a CLI-based environment.
State.
Figure 15: Lifecycle state transitions of the LinuxAgent.
The LinuxAgent’s lifecycle is governed by a minimal set of states, capturing the essential execution progression while maintaining simplicity and predictability:
  • CONTINUE: The task is ongoing and requires further CLI command execution to progress toward completion.
  • FAIL: The task cannot proceed under current system conditions or resources, signaling an unrecoverable error.
  • FINISH: The task has successfully completed all required operations.
At each step, the agent evaluates execution outcomes and determines the appropriate state transition, allowing the FSM to drive both normal progress and error handling in a structured, deterministic manner.
Strategy.
When in the CONTINUE state, the LinuxAgent executes a processor that orchestrates a small, modular set of strategies:
  • LLM_INTERACTION: Construct prompts using prior execution results and predefined templates to request next-step commands from the LLM.
  • ACTION_EXECUTION: Execute the CLI commands returned by the LLM, ensuring results are captured and structured for downstream processing.
  • MEMORY_UPDATE: Persist execution results and issued commands into the agent’s memory for future reference, enabling iterative refinement and error recovery.
This layered strategy design separates decision-making, such as prompt construction, from execution and state updates, enhancing modularity, reproducibility, and extensibility for future workflow modifications. In particular, unlike traditional polling-based or externally triggered data collection, the agent can proactively obtain system and environment information by invoking CLI commands on demand, eliminating unnecessary overhead and increasing responsiveness.
Command.
The LinuxAgent interacts with the MCP server via two primary commands:
  • EXEC_CLI: Execute arbitrary shell commands, capturing stdout and stderr for structured feedback.
  • SYS_INFO: Collect system-level information, such as memory usage, disk space, and hardware configuration, to inform decision logic or precondition checks.
These commands provide the atomic building blocks for the agent’s strategies, isolating system-specific operations within the client layer while enabling the server layer to focus on workflow orchestration and LLM-guided reasoning.
The LinuxAgent illustrates the minimal viable design for a single-agent system that integrates FSM control, strategy orchestration, and command execution. By maintaining a small, deterministic state set, modular strategies, and well-defined commands, the agent achieves robust, flexible, and traceable CLI task execution.
8.6.2 WindowsAgent: Multi-Agent Coordination on GUI Systems
Figure 16: Overall architecture of the WindowsAgent built upon UFO2. Figure adapted from the original paper.
While the LinuxAgent represents a lightweight, single-agent model for command-line environments, the WindowsAgent embodies a more sophisticated multi-agent framework tailored for GUI-based systems. We leverage the Desktop AgentOS UFO2 as the implementation of the WindowsAgent, which follows the same architectural principles introduced above but extends them to support multi-application coordination. Specifically, UFO2 consists of a HostAgent that decomposes a user request into multiple subtasks and assigns each subtask to an AppAgent, which executes the assigned subtask within an individual application. Below, we highlight the core design ideas and the major differences from LinuxAgent, and refer readers to the UFO2 paper for full details.
State.
The HostAgent adopts a state machine similar to that of the LinuxAgent, but introduces an additional ASSIGN state responsible for delegating subtasks to AppAgents. Once assigned, control transitions to the selected AppAgent for execution. Importantly, when an AppAgent reaches the FINISH or FAIL state, the control does not terminate but instead returns to the HostAgent, enabling it to decide whether to retry, re-plan, or advance to the next subtask. This hierarchical state transition mechanism naturally supports cooperative task completion across multiple applications.
Strategy.
Unlike LinuxAgent, where system states can be dynamically queried through CLI commands, GUI-based environments require explicit perception of the screen and interface hierarchy. Both HostAgent and AppAgent therefore begin each round with a DATA_COLLECTION strategy that captures screenshots and accessibility (a11y) metadata to construct a structured view of the GUI environment. These inputs are crucial for grounding subsequent reasoning and action decisions. The remaining strategies, LLM_INTERACTION, ACTION_EXECUTION, and MEMORY_UPDATE, follow the same modular workflow as in the LinuxAgent, and are omitted here for brevity. This design unifies the overall agent logic across heterogeneous platforms while allowing for system-specific customization.
Command.
The WindowsAgent exposes a significantly richer command set compared to the LinuxAgent. The HostAgent can invoke MCP tools to launch or select applications, while the AppAgent operates through a dedicated GUI MCP server capable of simulating user interactions such as mouse clicks and keyboard inputs. Furthermore, each AppAgent can integrate application-specific MCP servers that bridge to internal APIs, enabling faster and more reliable automation than purely vision-based approaches. This hybrid interaction model, combining GUI manipulation with API-level control strikes a practical balance between generality and robustness.
Overall, LinuxAgent and WindowsAgent jointly demonstrate how the proposed architecture can be adapted for both single-agent and multi-agent environments, from lightweight CLI systems to complex desktop ecosystems. The same design paradigm can be extended to other platforms, such as mobile or in-vehicle infotainment systems, ensuring scalability and reusability across diverse device types.
9 Implementation and Engineering
We implemented UFO3 as a large-scale system consisting of approximately core 73K lines of Python code, integrating the centralized ConstellationAgent, the asynchronous Constellation Orchestrator, the AIP communication layer, and 3 representative device agents, i.e.,, LinuxAgent, WindowsAgent and MobileAgent (Android). An additional 6.1K lines of code were developed for the frontend web UI. The implementation is further accompanied by over 77K lines of user documentation, detailing module interfaces, orchestration protocols, and configuration schemas. This extensive documentation ensures maintainability, reproducibility, and smooth multi-team integration, reflecting substantial engineering effort and demonstrating UFO3’s readiness for real-world deployment.
We leverage Python’s asyncio framework for concurrent event handling, allowing dynamic DAG updates, agent registration, and task execution to proceed asynchronously without blocking the orchestrator’s control loop. The system runs seamlessly across Linux and Windows environments, enabling dynamic workflow orchestration, persistent cross-device communication, and plugin-based extensibility.
9.1 Interactive UFO3 WebUI
Figure 17: Snapshot of the UFO3 WebUI. The interface integrates natural-language interaction, TaskConstellation visualization, and device agent management in real time.
We build a modern, futuristic WebUI that serves as the operator-facing control surface of UFO3, integrating chat interaction, real-time task monitoring, and device management within a single view (Figure 17). The chatbox at the center allows users to issue natural-language requests and observe agent reasoning and replies in real time. The right panel visualizes the evolving TaskConstellation as a DAG and lists all TaskStars with status indicators and detailed logs accessible via expansion. The lower-left registry displays all connected device agents with their capabilities, and states, enabling operators to quickly inspect, reconnect, or migrate tasks as needed. Execution events stream continuously through an event-driven backend built on FastAPI and WebSocket, ensuring sub-millisecond update latency and seamless synchronization between orchestrator and visualization.
This design provides high observability and transparency across the multi-agent workflow. Users can trace each decision from natural-language reasoning to execution outcomes, diagnose failures via per-TaskStar logs, and visualize dependency satisfaction in real time. By unifying conversation, orchestration, and monitoring, the WebUI bridges human intent and agentic execution, allowing rapid debugging, safe intervention, and fine-grained control without disrupting asynchronous task execution. Overall, the WebUI turns UFO3 from a background automation engine into an interactive, transparent, and trustworthy orchestration environment.
9.2 Plugin and Extension Framework
Both the ConstellationAgent and device agents expose configurable MCP servers interfaces implemented using the FastMCP package. Upon startup, each agent launches an embedded MCP server whose toolset is dynamically registered according to its role (e.g., Linux CLI tools, Windows GUI automation, or system telemetry collectors). This plugin mechanism allows developers to add new capabilities, such as a novel GUI driver or API connector, by implementing a lightweight interface, without altering any orchestration or protocol logic. This design significantly improves maintainability, reduces coupling, and enables rapid onboarding of new device types into the constellation.
9.3 Prompt and Model Integration
Prompts are modularly defined via a hierarchical configuration. Each agent maintains a core system prompt template augmented by a collection of in-context exemplars for few-shot adaptation. The ConstellationAgent centrally manages the LLM backend through a unified API layer compatible with OpenAI-style interfaces, enabling model-agnostic deployment across GPT-based, Claude-based, or local open-weight models. This separation between orchestration and inference ensures the entire system remains robust against future LLM model changes.
Figure 18: Example Markdown log generated by UFO3, showing agent actions, reasoning, TaskConstellation DAG (before and after edits).
9.4 Automated Task Logging.
To facilitate debugging and system introspection, UFO3 automatically generates detailed logs in Markdown format after each task execution. These logs capture the complete trace of agent actions, reasoning steps, and intermediate outputs, including thoughts, invoked commands, and returned results. They also include visualizations of the TaskConstellation DAG, showing both the initial and modified topologies, as well as task execution timelines (TaskStar and TaskStarLine). Figure 18 illustrates an example log output, demonstrating how these comprehensive records provide a clear, structured view of task execution for analysis and debugging.
10 Experimental Evaluation
Following the system implementation, we evaluate the UFO3 framework to understand its effectiveness, robustness, and coordination efficiency in realistic, heterogeneous environments. Existing agent benchmarks predominantly focus on single-device or single-OS settings (e.g., text-based API workflows or GUI automation) mu2025gui, which fail to capture the cross-device, cross-platform orchestration that UFO3 is designed for. Moreover, to the best of our knowledge, there exists no prior agent system capable of orchestrating multi-device tasks that span both Linux and Windows environments with unified control and shared context. This makes direct comparison against existing systems infeasible and potentially misleading.
Therefore, instead of benchmarking against prior single-agent or single-platform baselines, we focus on a comprehensive internal evaluation that characterizes UFO3’s performance across multiple dimensions, covering its planning accuracy, execution reliability, coordination efficiency, and fault tolerance. This approach allows us to isolate the impact of UFO3’s architectural innovations and evaluate how well it scales to realistic multi-device orchestration scenarios.
Our evaluation aims to answer the following research questions:
  • RQ1 (Task Completion): Can UFO3 successfully complete diverse, multi-agent tasks across heterogeneous devices and platforms?
  • RQ2 (Orchestration and Adaptation): How effectively does UFO3 orchestrate a user query into a structured TaskConstellation DAG, and how does it adapt the DAG dynamically in response to intermediate subtask results during execution?
  • RQ3 (Parallelism Exploitation): To what extent can UFO3 identify and exploit parallelism across independent subtasks to accelerate overall execution without compromising correctness?
  • RQ4 (Performance and Scalability): How efficient is UFO3 in task planning, scheduling, and end-to-end execution latency under different network conditions and system scales?
  • RQ5 (Robustness): How does UFO3 handle partial failures, network delays, or unavailable agents during distributed execution?
10.1 Experiment Setup
We deploy UFO3 across five physical machines in a controlled environment:
  • 1× Windows 11 desktop, running the WindowsAgent;
  • 3× Ubuntu 22.04 workstations (CPU-only), running LinuxAgent;
  • 1× Ubuntu 24.04 GPU node equipped with four NVIDIA A100 GPUs, running LinuxAgent.
Agents communicate via a simulated local network with 1–10 ms latency and a wide-area link (50–100 ms latency) for the GPU node using the AIP. All components are implemented in Python 3.10 and powered by the same large language model (GPT-5-Chat-20251003) openai2025gpt5systemcard for both the ConstellationAgent and device agents. The orchestrator and controller run on a dedicated management node that coordinates agent discovery, task scheduling, and monitoring. This setup emulates a realistic hybrid enterprise environment that combines cloud servers, local desktops, and GPU compute nodes.
10.2 NebulaBench: A Crossed-Device Evaluation Benchmark
Table 2: Overview of the 10 task categories used in evaluation. Each category includes 4–10 representative cases.
Category Description Count
Logs & Monitoring Log retrieval, aggregation, and report generation across devices. 6
System State & Configuration Managing environment variables, users, permissions, and disk information. 5
Processes & Services Starting, stopping, and monitoring services and scheduled tasks. 5
Data Wrangling & Scripting Parsing CSV/text files, performing statistical summaries, and executing scripts. 4
DevOps & Containers Managing Git repositories, CI/CD pipelines, and container operations. 10
Networking & Connectivity Diagnosing connectivity issues, pinging hosts, and verifying endpoints. 5
Browsing Tasks Web-assisted cross-device operations that require interacting with web resources via a browser and then applying or verifying results on remote Linux hosts. 5
Cross-Device Orchestration Coordinating multi-agent workflows involving data transfer and dependency handling. 5
GPU & ML Workloads Launching and monitoring GPU-based ML training and inference jobs. 5
Negative Tasks Infeasible or invalid tasks used to test failure detection and safe termination. 5
To assess UFO3’s generality and real-world applicability, we construct NebulaBench, a benchmark of 55 representative multi-agent tasks spanning typical productivity and system administration workflows. Five volunteers with diverse Linux and Windows experience proposed realistic queries they would naturally issue to a digital assistant in daily use. The resulting tasks span ten functional categories, summarized in Table 2. A detailed list of queries in NebulaBench is shown in Appendix C and Table LABEL:tab:galaxy_results.
Each scenario is labeled with one of three difficulty levels, namely Easy, Medium, and Hard, reflecting the degree of orchestration and reasoning required:
  • Easy: Single-host operations such as log inspection, one-off checks, service control, small file transfers, or short summarization. These tasks require minimal coordination and succeed with standard administrative privileges.
  • Medium: Moderately orchestrated tasks involving multiple machines or light cross-platform activity (e.g., container run-and-verify, metric aggregation, spreadsheet updates). They often require conditional logic, transient failure handling, or simple parsing and aggregation.
  • Hard: Complex multi-step workflows requiring cross-platform orchestration, CI/build pipelines, container image management, distributed data processing, or live patching. These tasks involve high privilege levels, non-trivial verification, and greater exposure to race conditions or dependency issues.
Figure 19 visualizes the composition of NebulaBench. The left pie chart shows the distribution of task difficulty, which is roughly balanced across Easy (33%), Medium (35%), and Hard (33%) tasks. The right chart depicts the number of devices required per task, ranging from 0 to 5. Tasks requiring no devices correspond to negative tasks that are inherently infeasible; the system is expected to detect and fail these safely without executing any agent actions. Most tasks, however, involve 3–5 devices, highlighting NebulaBench’s focus on multi-agent orchestration scenarios.
This stratification enables a controlled evaluation of UFO3’s reasoning, coordination, and execution robustness under increasing complexity. Each task is executed end-to-end, from natural language request to final result through the orchestrator–agent hierarchy, with all required data, scripts, and code pre-deployed. By including negative tasks, we also validate UFO3’s safety mechanisms, ensuring that infeasible goals are correctly identified and aborted without unintended side effects.
Figure 19: Distribution of NebulaBench tasks. Left: proportion of tasks by difficulty. Right: number of devices involved per task.
10.3 Metrics and Methodology
To rigorously assess the effectiveness and robustness of UFO3, we design a comprehensive, multi-dimensional evaluation methodology covering correctness, adaptivity, efficiency, and fault tolerance. Our goal is not only to measure whether tasks are completed, but also to understand how well UFO3 plans, adapts, and sustains performance under dynamic and heterogeneous conditions. We organize our analysis around four research questions (RQ1–RQ4), each corresponding to a key aspect of UFO3’s system behavior.
RQ1: Task Completion.
We measure the overall Task Success Rate (TSR) and Subtask Completion Rate (SCR) to evaluate UFO3’s execution reliability. TSR captures the fraction of user queries that are successfully completed end-to-end, as judged by the query authors based on final outcomes. SCR measures the success rate of individual subtasks executed by device agents, annotated automatically by an LLM evaluator that inspects the task trajectories and completion logs. Together, TSR and SCR provide a top-down and bottom-up view of system reliability across the orchestration hierarchy. Note that for negative test cases, we mark a task or subtask as successful when the agent correctly detects and reports the intended failure, rather than attempting to complete it erroneously.
RQ2: Orchestration and Adaptation.
We evaluate UFO3’s reasoning and dynamic re-planning by analyzing the initial TaskConstellation DAG and its evolution during execution. To quantify adaptability, we track DAG modification metrics, including edits, node insertions and deletions, and dependency changes, which reflect how the system adjusts the task constellation in response to runtime results. Comparing the initial and final DAGs provides a holistic measure of UFO3’s ability to adapt its plan as execution unfolds.
RQ3: Parallelism and Execution Efficiency.
To evaluate UFO3’s ability to exploit concurrency and optimize execution, we measure several metrics derived from the TaskConstellation DAG:
  • Maximum Parallel Width: the largest number of subtasks that can be executed concurrently at any point in the DAG.
  • Critical Path Length (L): the execution time of the longest serial dependency chain in the DAG.
  • Total Task Execution Time (W): the sum of execution times of all tasks, representing the cumulative workload.
  • Parallelism Ratio (P): $P=W/L$, capturing the degree to which UFO3 leverages parallel execution across devices.
These metrics provide insight into UFO3’s efficiency in orchestrating multi-agent tasks and its effectiveness in parallelizing independent subtasks to reduce overall task latency.
RQ4: Latency and Scalability.
We evaluate End-to-End Latency, Orchestration Time (including TaskConstellation creation and edits), and Total Task Execution Time to characterize system efficiency. These metrics capture the balance between reasoning overhead, arising from LLM-based planning, scheduling, and coordination, and the execution gains achieved through distributed parallelism. Together, they provide a holistic view of UFO3’s scalability across heterogeneous agents and devices under realistic workload conditions.
RQ5: Robustness and Fault Handling.
We assess robustness through a case study on a single request executed under three controlled failure modes: (i) a single device agent fails and recovers via dynamic DAG reassignment, (ii) a single agent fails without recovery, and (iii) all agents fail simultaneously. These scenarios capture UFO3’s fault tolerance boundary and illustrate its adaptive recovery mechanisms under partial and global disruptions.
Overall, this evaluation framework enables a holistic examination of UFO3’s behavior under diverse operational conditions, revealing how design principles such as event-driven scheduling, runtime DAG modification, and agent autonomy translate into measurable system benefits.
10.4 RQ1: Task Completion Analysis
Figure 20: Task and subtask success rates by difficulty level (left) and by query category (right) of UFO3.
To evaluate UFO3’s effectiveness in executing distributed tasks, we first examine the primary metrics of interest: Subtask Completion Rate (SCR) and Task Success Rate (TSR). We analyze performance both by task difficulty and by functional category, as summarized in Figure 20  [^1] A complete list of UFO3’s performance on all queries is provided in Appendix C and Table LABEL:tab:galaxy_results.. Overall, UFO3 demonstrates strong reliability across the benchmark, achieving an SCR of 83.3% and a TSR of 70.9%. These high rates indicate that the orchestrator effectively coordinates subtasks across devices, while individual device agents reliably execute their assigned operations, validating the robustness of UFO3’s design. Breaking down by difficulty, easy and medium tasks exhibit comparable TSRs (72.2% and 73.7%, respectively), showing that UFO3 handles routine and moderately interdependent tasks effectively. For hard tasks, TSR declines to 66.7%, yet even in these complex multi-host scenarios UFO3 successfully completes the majority of requests, highlighting its capability under challenging orchestration conditions.
By functional category, performance trends align with task characteristics. Structured, deterministic tasks such as Data (SCR 100%) and Proc (SCR 96%) achieve the highest reliability. Tasks involving user-like interactions or ambiguity, including Browsing (SCR 64.2%) and Negative scenarios (SCR 66.7%), show lower success rates. Intermediate categories requiring cross-device coordination, such as Orchestration (SCR 85%) and DevOps (SCR 77.1%), demonstrate solid subtask reliability, evidencing UFO3’s capacity to manage dependencies across heterogeneous devices.
Across both analyses, SCR generally exceeds TSR. This gap arises because a single subtask failure among multiple dependencies can cause an overall task failure, even when most subtasks succeed. The pattern underscores the value of UFO3’s dynamic DAG management and reasoning, which mitigate partial failures and maintain high system reliability in complex, distributed workflows.
Error Analysis.
To better understand the remaining failures, we examined common error patterns. First, tasks requiring file transfers across devices occasionally fail because AIP currently supports only textual communication; device agents may not have direct network access to each other. Future work includes maintaining a shared memory and extending AIP to support file read/write operations. Second, some failures arise when agents attempt to complete tasks regardless of preconditions. For example, the request “Start UFO3 service on all Linux devices” triggered the LinuxAgent to create the service even when it did not exist, rather than reporting a failure. Enhancing agent self-awareness and enforcing conditional execution will mitigate such issues. Third, tasks executed via WindowsAgent show relatively lower success due to GUI dependencies and interface variability; improving robustness of GUI-based agents remains an important direction.
Despite these errors, UFO3’s generated TaskConstellation are largely correct, indicating that the ConstellationAgent effectively decomposes complex requests, identifies dependencies, and exposes parallelism, validating its reasoning design.
10.5 RQ2: Orchestration and Adaptation
Table 3: Average modifications per edit and per request by type.
Added
Tasks
Removed
Tasks
Modified
Tasks
Added
Dependencies
Removed
Dependencies
Modified
Dependencies
Total
Per Edit 0.05 0.00 0.99 0.04 0.01 0.00 1.09
Per Request 0.41 0.00 5.20 0.24 0.07 0.00 5.91
Figure 21: Comparison of the initial and final task (left) and dependency (right) counts in the TaskConstellation of UFO3.
We next examine UFO3’s ability to orchestrate a user query into an executable TaskConstellation DAG and adapt it dynamically as execution unfolds. To quantify this, we track modifications to the DAG across all requests, including node insertions, deletions, and dependency updates in Figure 21, and compare the initial and final task and dependency counts in Table 3.
Overall, UFO3 demonstrates a highly active editing behavior, with an average of 1.09 modifications per edit and 5.91 modifications per request. Most changes occur in the Modified Tasks category, reflecting the system’s strategy of enriching downstream tasks based on the results of preceding subtasks. For example, logs collected by earlier tasks are incorporated into subsequent document-writing or analysis subtasks, ensuring that later steps have complete context and accurate information. In contrast, the number of added or removed tasks and dependencies is minimal, indicating that ConstellationAgent generally produces well-structured DAGs from the outset and only fine-tunes tasks during execution.
Breaking down by difficulty (Figure 21), we observe that UFO3 maintains a robust initial decomposition across Easy, Medium, and Hard tasks, with only modest increases in both task and dependency counts by the end of execution. This demonstrates that while the orchestrator adapts to runtime results, it does so without fundamentally restructuring the workflow, preserving overall plan stability. Harder tasks show slightly more edits, consistent with their longer multi-step workflows and richer context propagation.
Taken together, these observations indicate that UFO3 exhibits strong adaptive orchestration capabilities: (i) it actively enriches downstream tasks based on prior subtask outputs, ensuring context-aware execution; (ii) it generates high-quality initial DAGs, requiring minimal structural modifications; and (iii) it balances stability with runtime flexibility, applying refinements without disrupting the overall task plan. These features highlight UFO3’s ability to both plan effectively and adjust dynamically across distributed devices.
10.6 RQ3: Parallelism and Execution Efficiency
Table 4: Parallelism characteristics of task constellations by difficulty.
Difficulty
Max
Parallel Width
Critical Path
Length (L)
Total Exec.
Time (W)
Parallelism
Ratio (P = W/L)
Easy 3.53 144.10 315.28 1.86
Medium 2.73 166.94 212.54 1.51
Hard 3.21 232.12 339.65 1.77
Overall 3.17 178.34 289.20 1.72
We next investigate UFO3’s ability to exploit concurrency and optimize distributed task execution. Table 4 reports four DAG-derived metrics that capture the structural and runtime efficiency of each task constellation. Across all scenarios, UFO3 achieves an average parallelism ratio of 1.72, with up to 3.5 concurrent subtasks executing at peak, demonstrating that UFO3 can effectively uncover and schedule independent subtasks across devices.
We observe several trends. First, the maximum parallel width remains consistently high (around 3 tasks) even as task difficulty increases, indicating that UFO3’s orchestrator maintains concurrency opportunities even in complex, multi-host settings. Second, while the critical path length naturally grows with task complexity (144s $\to$ 232s), the increase in total execution time (315s $\to$ 340s) is modest, showing that UFO3 successfully overlaps execution through concurrent scheduling. Third, medium-difficulty tasks exhibit slightly lower parallelism ($P=1.51$), as many involve short verification or data aggregation steps with limited parallel components, while both easy and hard tasks benefit more from concurrent flows.
These results reveal that UFO3’s DAG-based orchestration is not only structurally well-parallelized but also runtime-efficient. By modeling tasks as DAGs, UFO3 naturally identifies and executes independent subtasks asynchronously across heterogeneous devices, keeping the system utilization high. Together, these findings demonstrate that UFO3 delivers strong execution efficiency through fine-grained parallel orchestration, validating the effectiveness of its DAG-based design.
Figure 22: Task timing and orchestration breakdown by difficulty.
Figure 23: Fault injection scenarios for robustness evaluation. UFO3’s adaptive behavior under (1) transient, (2) permanent, and (3) full-agent failures.
10.7 RQ4: Latency and Scalability
We evaluate End-to-End Latency, Orchestration Time (including TaskConstellation creation and edits), and Total Task Execution Time to characterize UFO3’s efficiency and scalability. These metrics reflect the trade-off between reasoning overhead, arising from LLM-based planning, scheduling, and coordination, and execution gains obtained through distributed parallelism.
As shown in Figure 22 (left), UFO3 exhibits strong execution efficiency across all task difficulties. On average, the measured end-to-end latency is only 243.7 seconds, notably shorter than the combined orchestration and task execution time of 352.5 seconds (63.3 + 289.2). This corresponds to a 31% reduction in total completion time compared to a fully sequential workflow. Even for complex, cross-device tasks, UFO3 maintains an average end-to-end latency of about 4 minutes, demonstrating its ability to exploit distributed parallelism while preserving robust coordination and correctness. Compared to manual execution, which often takes tens of minutes or longer, UFO3 delivers a substantial improvement in both speed and scalability. This efficiency gain stems from UFO3’s concurrent design described in Section 6.2: multiple subtasks are executed in parallel across heterogeneous agents, while ConstellationAgent continues to refine the TaskConstellation asynchronously in the background. Such overlap between orchestration and execution effectively hides reasoning latency and maximizes device utilization.
We also observe a clear correlation between task difficulty and latency: harder tasks incur longer orchestration times (84.2s for Hard vs. 55.1s for Easy) and longer end-to-end latency (325.2s vs. 195.9s), reflecting the increased complexity, number of subtasks, and cross-device coordination required. However, even for Hard tasks, the total end-to-end time remains within a few minutes, demonstrating UFO3’s efficiency compared to manual execution of equivalent multi-device workflows.
Figure 22 (right) further breaks down the average time spent in TaskConstellation creation, editing, and individual task completion. Creation and editing are consistently fast (16–20s and 8–12s, respectively), indicating that ConstellationAgent is highly efficient at planning and adapting task DAGs in real time. Task completion, averaging around 48s overall, dominates the workflow, yet remains reasonable given the distributed nature and heterogeneity of devices.
These results underscore several key insights: (i) UFO3 leverages cross-device parallelism effectively, enabling task execution to substantially overlap with reasoning and orchestration; (ii) the orchestration engine scales gracefully with task complexity, keeping latency manageable even for multi-step, multi-host tasks; (iii) the efficient creation and editing of the TaskConstellation demonstrates that ConstellationAgent can dynamically adapt plans with minimal overhead, supporting robust and responsive execution in real-world scenarios. Collectively, these observations confirm that UFO3 is both performant and scalable for heterogeneous, distributed multi-agent workloads.
10.8 RQ5: Robustness and Fault Handling
Finally, we evaluate UFO3’s robustness and fault-tolerance capabilities by observing its behavior under three simulated device-agent failure scenarios. The test request is: “Run long_job.sh concurrently on Linux 1–3 and report their running time on Notepad.” Ideally, this request invokes three LinuxAgent instances to execute the script in parallel, followed by the WindowsAgent aggregating and recording the results in Notepad. We design three fault-injection scenarios to examine UFO3’s adaptive responses:
  • Scenario 1: One LinuxAgent is intentionally disconnected but recovers shortly before the task completes.
  • Scenario 2: The same LinuxAgent disconnects and does not recover for the remainder of the execution.
  • Scenario 3: All LinuxAgent instances remain unavailable until the task ends.
These cases, illustrated in Figure 23, allow us to analyze how UFO3 detects failures, adapts the TaskConstellation accordingly, and maintains graceful degradation or recovery during runtime.
In Scenario 1, where Linux 1 is temporarily disconnected, UFO3 detects the failure event immediately. The corresponding subtask (Task A) is marked as failed, and upon receiving this signal, the ConstellationAgent proactively spawns a replacement task (Task A’) to retry execution. When Linux 1 reconnects before the task deadline, the retry is successfully executed, and the results are correctly aggregated by the WindowsAgent. This demonstrates UFO3’s ability to automatically recover from transient device failures through adaptive re-planning without human intervention, showing that its fault-handling mechanism is both responsive and self-healing.
In Scenario 2, Linux 1 remains disconnected throughout execution. UFO3 follows the same retry logic, creating a new Task A’ and reassigning it to the same device after an adaptive delay. However, as the device remains unavailable, the retry also fails. Instead of hanging indefinitely or aborting the entire workflow, UFO3 recognizes this as a partial failure, continues executing the remaining subtasks on other available devices, and finally records the aggregated results, including explicit failure traces in the final Notepad report. This behavior shows that UFO3’s orchestration framework can degrade gracefully under partial failures, preserving useful progress and ensuring result completeness.
In Scenario 3, all LinuxAgent instances are disconnected, leading to consecutive retry failures across all subtasks. After exhausting retry attempts, the ConstellationAgent terminates the execution plan and reports the entire request as failed. The failure is propagated in a transparent manner, and UFO3 refrains from producing hallucinated or incomplete results. This highlights the system’s emphasis on fail-safe integrity, where it prefers honest termination over speculative completion when recovery is infeasible.
Overall, these experiments demonstrate that UFO3 exhibits a high degree of robustness in distributed environments. Its adaptive orchestration loop allows it to distinguish between transient and permanent failures, retry safely, and preserve progress when possible. More importantly, the system maintains end-to-end consistency even under multi-agent disruptions, validating the effectiveness of its hierarchical fault-tolerance design.
10.9 Case Study: Cross-Device Orchestration in Action
Lastly, we present three representative cases from NebulaBench to demonstrate how UFO3 effectively and efficiently orchestrates heterogeneous device agents to complete complex user requests across operating systems and application boundaries.
10.9.1 Case 1: Collecting Logs and Writing Reports
Figure 24: Case 1: Log collection and report generation. UFO3 decomposes the high-level request into parallel log retrieval tasks on Linux agents and sequential report-generation tasks on Windows, demonstrating cross-device coordination and result aggregation.
As illustrated in Figure 24, this case requests UFO3 to “Retrieve all warning and error logs from Linux servers, add them to the ‘report’ sheet of the log_detailed.xlsx file, and send an email with the report to the operations engineer.” While conceptually straightforward, this task traditionally requires tedious manual effort across multiple systems and interfaces.
UFO3 autonomously decomposes the request into five TaskStars: three for parallel log retrieval and two for final report generation and email dispatch. The first three TaskStars are dispatched to distinct LinuxAgents, each collecting logs asynchronously from its assigned host. Once completed, the results are automatically aggregated by the ConstellationAgent, which then spawns two sequential TaskStars on the WindowsAgent to insert the logs into Excel and compose an email summary.
This case exemplifies UFO3’s strong parallelism and adaptive coordination capabilities. By overlapping asynchronous data collection with centralized result fusion, UFO3 achieves both time efficiency and consistency across heterogeneous systems, significantly reducing human effort and turnaround time compared to manual workflows.
10.9.2 Case 2: Task Allocation and Result Integration via Excel
Figure 25: Case 2: Distributed task execution from a centralized Excel sheet. UFO3 interprets structured task descriptions from Excel, distributes them to Linux agents for execution, and writes concise summaries back into the corresponding cells.
In Figure 25, we illustrate the second case, where UFO3 is asked to “Complete all tasks listed in schedule.xlsx on Windows using Linux servers, and write back a one-sentence result summary to the Result Summary column.” This case evaluates UFO3’s ability to bridge structured spreadsheet data manipulation with distributed execution across heterogeneous devices.
Upon receiving the request, UFO3 first creates a TaskStar on the WindowsAgent to parse the Excel sheet, identifying five distinct subtasks, each targeting a specific Linux host. These subtasks (Task B–D) are then dispatched to the respective LinuxAgents for concurrent execution. After all remote tasks complete, the ConstellationAgent aggregates their outputs, synthesizes concise summaries, and writes them back into the appropriate cells of the same Excel file.
This case demonstrates how UFO3 seamlessly integrates centralized data management with distributed execution, effectively turning Excel into a live orchestration interface. It highlights the system’s fine-grained understanding of data provenance, automatically preserving task-to-cell correspondence without human supervision.
10.9.3 Case 3: Resource-Aware Distributed Computation
Figure 26: Case 3: Resource-aware orchestration of distributed computation. UFO3 dynamically inspects hardware and workload conditions on four Linux agents, assigns tasks based on available resources, and generates a human-readable summary report on Windows.
Finally, Figure 26 presents a more heterogeneous and complex example, where the user requests: “I have a distributed computation task that requires coordinating four Linux hosts. The task consists of GPU-intensive matrix multiplication, CPU-intensive dataset processing, and memory-intensive data aggregation. Please inspect hardware and current workload on all four hosts, then assign one GPU-intensive task, two CPU-intensive tasks, and one memory-intensive task to the appropriate hosts. Provide a task assignment report including host specs, assigned workloads, and notes in Notepad on Windows.”
UFO3 first triggers all four LinuxAgents to perform hardware inspection and runtime workload evaluation in parallel (Task A–D). Based on the collected telemetry, the ConstellationAgent dynamically plans the task distribution strategy, ensuring that computationally expensive workloads are placed on the most suitable hosts. After the distributed executions complete, the WindowsAgent automatically compiles a task assignment report summarizing host capabilities, workload mapping, and execution notes in a human-readable format.
This case highlights UFO3’s capacity for adaptive, resource-aware orchestration, combining multi-device reasoning, real-time environment sensing, and distributed execution under a unified agentic framework. The entire workflow, from information gathering to decision-making and reporting, is autonomously coordinated, underscoring UFO3’s ability to generalize beyond predefined templates and operate effectively in dynamic, real-world computing scenarios.
Summary.
Across these three representative cases, UFO3 consistently demonstrates strengths in parallelization, cross-application integration, and adaptive reasoning. It effectively decomposes high-level user goals into executable subtasks, dispatches them across heterogeneous devices, and synthesizes coherent outputs in natural language or structured files. Together, these results illustrate UFO3’s promise as a general-purpose, LLM-powered orchestration system capable of automating diverse, realistic workflows with high efficiency, transparency, and reliability.
11 Discussion
11.1 Scalability and Ecosystem Growth
Although the current UFO3 prototype integrates Windows and Linux device agents, the underlying architecture is inherently platform-agnostic. The AIP and device agent template enable rapid extension to new platforms such as mobile, tablet, or IoT devices. Developers can implement lightweight clients that conform to the AIP specification and seamlessly connect to the orchestrator without modifying the core system.
Moreover, UFO3 orchestrates at the protocol level rather than binding to specific implementations. Any agent compatible with AIP, whether it is a coding agent yang2024swe, a deep-research agent huang2025deep, or a data-analysis module zhang2025allhands, can participate in the orchestration. This design opens the door for a broader agent ecosystem, where specialized agents cooperate under a unified orchestration fabric. Our long-term vision is to make UFO3 an “agent galaxy”, a shared runtime that interconnects diverse autonomous agents into a cohesive, cross-platform execution universe.
11.2 Limitations of Device-Centric Execution
At the core of UFO3 lies the ConstellationAgent, which governs dynamic DAG scheduling and reasoning. However, the system’s overall capability remains constrained by the competence of individual device agents. Since execution is distributed, a single agent’s failure, such as incomplete subtask execution or incorrect result reporting, can cascade through the constellation and lead to global task failure cemri2025multi. This “weakest-link effect” underscores the need for more reliable and self-verifying device agents. Future work will explore automatic capability validation, task-level sandboxing, and adaptive reallocation strategies that allow the orchestrator to detect and recover from unreliable nodes without compromising the consistency of the global task state.
11.3 Toward Shared Cross-Device Memory
Currently, agent communication in UFO3 is purely textual, suitable for command exchange but insufficient for data-rich workflows involving images, logs, or binaries. Many cross-device tasks, however, require file transfer or shared context, for example, a Windows agent generating a dataset for a Linux node to process. To support such workflows, we plan to extend AIP with a persistent, shared “agent memory” zhang2025survey, effectively a distributed workspace where agents can store and retrieve intermediate results. This enhancement would unify context across devices, simplify file sharing, and enable richer collaborative behavior across heterogeneous agents.
12 Related Work
We review related research on digital device agents and multi-agent orchestration systems.
12.1 Digital Device Agents
The emergence of large language models, particularly multimodal variants, has enabled a new class of intelligent agents capable of operating directly on digital devices zhang2024large. These agents extend beyond web and mobile environments to desktops, and are now expanding into domains such as automotive and embedded systems.
Early systems like SeeAct zheng2024gpt, Mobile-Agent wang2024mobile, and UFO zhang2025ufo pioneered this direction by leveraging GPT-4V yang2023dawn to perceive graphical user interfaces and perform human-like interactions across different platforms. Subsequent research has evolved along two main trajectories, namely (i) system-level integration, enhancing robustness through tighter coupling with native OS and API layers zhang2025api; zhang2025ufo2; and (ii) model-level adaptation, fine-tuning LLMs with large-scale interaction data and reinforcement learning to improve grounding and autonomy in digital environments luo2025gui; zheng2025vem; wang2025ui.
Industry adoption has accelerated rapidly. Anthropic’s Claude (Computer Use) anthropic2024 relies purely on screenshot-based perception, while OpenAI’s Operator cua2025 demonstrates strong multimodal reasoning and robust desktop automation. These advances collectively signal the formation of a new computing paradigm where LLMs serve as universal control interfaces for heterogeneous software systems.
Despite this progress, most existing device agents remain confined to isolated machines, limiting context sharing and collaborative capability. UFO3 addresses this limitation by introducing cross-device orchestration. Through the AIP, UFO3 unifies disparate agents into a coherent constellation, enabling seamless coordination, shared state, and large-scale cooperative intelligence across devices.
12.2 Multi-Agent Orchestration
Recent years have witnessed growing interest in LLM-based multi-agent systems, where coordination among heterogeneous agents becomes central to achieving emergent intelligence guo2024large. Designing robust orchestration protocols, covering role assignment, communication, collaboration, and context sharing, has been identified as a key challenge bhatt2025should; dang2025multi; kong2025survey.
The Internet of Agents (IoA) cheninternet represents one of the earliest structured attempts in this direction. It introduces an integration protocol and an instant-messaging-style communication layer, enabling flexible, scalable multi-agent collaboration. IoA emphasizes dynamic teaming and conversation flow control, allowing agents to self-organize and cooperate fluidly in an Internet-like environment yang2025agentic. Building upon this idea, the Federation of Agents (FoA) giusti2025federation further advances distributed orchestration with Versioned Capability Vectors (VCVs) as machine-readable profiles that encode agent capabilities, costs, and constraints. These semantic embeddings make capabilities discoverable and composable, enabling dynamic, capability-driven coordination at scale.
In contrast, UFO3 extends this vision into the realm of heterogeneous digital devices. It provides a unified, extensible orchestration substrate that connects agents distributed across desktops, browsers, and operating systems. Through the AIP, UFO3 not only facilitates low-latency cross-device communication and shared context propagation but also enables composable construction of agent teams. This design transforms isolated device agents into a coherent digital ecosystem, scalable, evolvable, and self-organizing.
13 Conclusion
We presented UFO3, a cross-device orchestration system that transforms LLM agents from isolated executors into coordinated collaborators. At its core, UFO3 introduces the mutable TaskConstellation abstraction, which tightly integrates planning and execution: the ConstellationAgent continuously synthesizes and edits DAGs; the Constellation Orchestrator schedules tasks asynchronously and applies safe, batched updates while enforcing invariants that guarantee single assignment and DAG acyclicity; and the Agent Interaction Protocol (AIP) provides persistent, low-latency, and resilient communication across heterogeneous devices. A layered device-agent template, backed by MCP servers, allows new endpoints to join seamlessly without modifying the control plane.
Our evaluation on NebulaBench, comprising 55 representative tasks across 5 machines and 10 categories, demonstrates that UFO3 reliably orchestrates heterogeneous workflows: achieving 83.3% subtask completion, 70.9% task success, an average parallelism ratio of 1.72, and 243.7 s mean end-to-end latency, 31% faster than a sequential baseline thanks to effective orchestration–execution overlap. Fault-injection experiments further confirm UFO3’s robustness: it gracefully recovers from transient agent failures, preserves partial progress under persistent outages, and conservatively terminates under global failures.
Overall, UFO3 lays the foundation for the next generation of autonomous multi-agent systems, enabling rich memory sharing, seamless coordination, and expansive device integration. By weaving heterogeneous devices and intelligent agents into a unified fabric, UFO3 brings us closer to an adaptive and reliable ecosystem of digital assistants that act collectively as a super-agent to harness ubiquitous intelligence.
Appendix A Formal Guarantees and Model
State and Transitions.
We model the orchestrator as an asynchronous transition system with state
$\sigma\;=\;(C,\,s,\,A,\,L,\,Q,\,D),$
where $C=(V,E)$ is the task-constellation DAG over a finite task universe $T$ ($V\!\subseteq\!T$, $E\!\subseteq\!V\times V$); $s:V\to\{\textsf{PENDING},\textsf{RUNNING},\textsf{COMPLETED},\textsf{FAILED}\}$ is the task-state map; $A:V\rightharpoonup\mathcal{D}$ is the partial device assignment over devices $\mathcal{D}$; $L\in\{\textsf{free},\textsf{held}\}$ is the global edit/assignment lock; $Q\in\operatorname{Seq}(\mathcal{E})$ is the pending event queue over $\mathcal{E}=\{\langle\textsf{TASK\_COMPLETED},t\rangle,\langle\textsf{TASK\_FAILED},t\rangle\mid t\in V\}$; and $D\subseteq\mathcal{D}$ is the set of available devices. A task $t$ is ready in $(C,s)$ iff
$\textsf{ready}_{C}(s,t)\;\equiv\;s(t)=\textsf{PENDING}\ \wedge\ \forall(u,t)\!\in\!E.\ s(u)=\textsf{COMPLETED}.$
The system alternates between a lock-bounded editing phase (apply $\Delta$, validate, synchronize, publish) and an unlocked scheduling phase.
Inductive Invariants.
  • I1 (Single assignment during run). $A$ is a partial function $V\rightharpoonup\mathcal{D}$; if $s(t)=\textsf{RUNNING}$ then $A(t)\in D$ and $A(t)$ does not change while $t$ remains RUNNING.
  • I2 (Acyclic consistency). $C$ remains a DAG after every committed edit.
  • I3 (Edit locality / immutability). Within one edit cycle, edits may modify only PENDING tasks and edges whose endpoints are PENDING; RUNNING/COMPLETED/FAILED nodes and their incident edges are immutable.
Safety.
Each edit–validate–synchronize–publish–release cycle is linearized as a single atomic operation with a linearization point between publishing the change and releasing the lock. As dispatch occurs only when $L=\textsf{free}$, all assignments are taken w.r.t. a validated DAG. A standard induction over the transition relation shows that I1–I3 are preserved.
Liveness and Deadlock Freedom.
Let $R=\{t\in V\mid s(t)=\textsf{RUNNING}\}$ at lock acquisition. While $L=\textsf{held}$, no new RUNNING tasks are created, so only completions/failures from $R$ can arrive. With the variant $\Phi(\sigma)=|Q|+|R|$, every sync or completion event decreases $\Phi$; hence the edit phase terminates and the lock is released. Under weak fairness of event delivery and device availability, every perpetually ready task is eventually dispatched.
Edit–Sync Confluence.
Let $\mathrm{Apply}(C,\Delta)$ be an edit that respects I3 and preserves acyclicity, and let $\mathrm{Sync}(s,E)$ fold a multiset $E$ of completion/failure events into $s$ via a per-task monotone, idempotent join. Because $\mathrm{Apply}$ leaves $s$ unchanged and $\mathrm{Sync}$ leaves $C$ unchanged, and they operate on disjoint components (I3), the two maps commute within the same lock interval; the observable post-state $(C^{\prime},s^{\prime})$ is independent of whether events are folded before or after $\Delta$.
A.1 TLA++ Specification
Scope. The spec below matches the model above. For model-checking small instances, we stub Acyclic as TRUE and Apply/Synchronize as no-ops; safety is still meaningfully exercised (I1, I2), and the environment is bounded by a queue-length predicate QueueBound. Fairness assumptions are embedded in Spec.
---- MODULE Orchestrator ----
EXTENDS Naturals, Sequences
CONSTANTS TASKS, DEVICES
TaskStates == {"PENDING", "RUNNING", "COMPLETED", "FAILED"}
TaskEvents == {"TASK_COMPLETED", "TASK_FAILED"}
VARIABLES C, S, A, L, Q, D
QLen == Len(Q)
QueueBound == QLen <= 2
NULL == "NULL"
Acyclic(g) == TRUE
IsDAG(C_) == Acyclic(C_)
Ready(t) ==
/\ t $\in$ C.V
/\ S[t] = "PENDING"
[…]
TypeOK ==
/\ C $\in$ [V : SUBSET TASKS, E : SUBSET ($\X TASKS$)]
/\ S $\in$ [TASKS -> TaskStates]
/\ A $\in$ [TASKS -> (DEVICES $\cup$ {NULL})]
/\ L $\in$ {"free", "held"}
/\ Q $\in$ Seq(TaskEvents)
/\ D $\subseteq$ DEVICES
[…]
I2 == IsDAG(C)
LockFree == (L = "free")
LockHeld == (L = "held")
Init ==
/\ C = [V |-> TASKS, E |-> {}]
/\ S = [t $\in$ TASKS |-> "PENDING"]
/\ A = [t $\in$ TASKS |-> NULL]
/\ L = "free"
/\ Q = << >>
/\ D = DEVICES
/\ TypeOK
Enqueue ==
[…]
/\ Q = Append(Q, e)
/\ UNCHANGED <<C, S, A, L, D>>
Acquire ==
/\ LockFree
/\ L = "held"
/\ UNCHANGED <<C, S, A, Q, D>>
Release ==
/\ LockHeld
/\ L = "free"
/\ UNCHANGED <<C, S, A, Q, D>>
Edges(C_, t) == { u $\in$ C_.V : (<<u, t>> $\in$ C_.E) \/ (<<t, u>> $\in$ C_.E) }
Apply(C_, Delta, e) == C_
Synchronize(S_, C1) == S_
EditStep ==
/\ LockHeld
/\ Len(Q) > 0
/\ LET e == Head(Q) IN
LET Q1 == Tail(Q) IN
[…]
C1 $\in$ [V : SUBSET TASKS, E : SUBSET ($\X TASKS$)],
S1 $\in$ [TASKS -> TaskStates] :
/\ C1 = Apply(C, Delta, e)
/\ IsDAG(C1)
/\ S1 = Synchronize(S, C1)
[…]
(S[t] $\in$ {"RUNNING","COMPLETED","FAILED"} =>
/\ S1[t] = S[t]
/\ Edges(C1, t) = Edges(C, t))
/\ C = C1
/\ S = S1
/\ Q = Q1
/\ UNCHANGED <<A, L, D>>
DrainOrNoop ==
\/ (LockHeld /\ Len(Q) = 0 /\ UNCHANGED <<C, S, A, L, Q, D>>)
\/ EditStep
Dispatch ==
[…]
/\ LockFree
/\ Ready(t)
/\ A[t] = NULL
/\ S = [S EXCEPT ![t] = "RUNNING"]
/\ A = [A EXCEPT ![t] = d]
/\ UNCHANGED <<C, L, Q, D>>
UpdateDevices ==
[…]
/\ D = Dnew
/\ UNCHANGED <<C, S, A, L, Q>>
Noop == UNCHANGED <<C, S, A, L, Q, D>>
Next ==
\/ Enqueue
\/ Acquire
\/ DrainOrNoop
\/ Release
\/ Dispatch
\/ UpdateDevices
\/ Noop
vars == <<C, S, A, L, Q, D>>
Spec ==
/\ Init
/\ [][Next]_vars
/\ WF_vars(Dispatch)
/\ WF_vars(DrainOrNoop)
/\ WF_vars(Release)
THEOREM Spec => [](TypeOK /\ I1 /\ I2)
A.2 Model-Checking Configuration and Results
Configuration.
We evaluate safety and deadlock-freedom on small instances with a bounded event window (§A). The queue bound is expressed in the specification via QueueBound; we constrain it in the model as follows.
-- Orchestrator.cfg
SPECIFICATION Spec
CONSTANTS
TASKS = {t0, t1, t2}
DEVICES = {"dev0", "dev1", "dev2"}
CONSTRAINT QueueBound
INVARIANTS
TypeOK
I1
I2
CHECK_DEADLOCK TRUE
Results.
TLC completes exploration under the above configuration with the following summary:
  • 93,63393{,}633 states generated; 7,1687{,}168 distinct states; queue empty at fixpoint.
  • Graph search depth: $8$; average outdegree: $1$ (min 0, max $19$, $95$th percentile $8$).
  • Action-level distinct states: Init $1$, Enqueue $6$, Acquire $448$, Dispatch $441$, UpdateDevices 6,2726{,}272; others 0.
  • Fingerprint collision probability: $3.4\times 10^{-11}$.
All invariants (TypeOK, I1, I2) hold and TLC reports no deadlocks.
Reproducibility notes.
Weak fairness on Dispatch, DrainOrNoop, and Release is enabled in Spec. The QueueBound predicate enforces the bounded-window assumption used in the liveness argument. For larger instances or unbounded event ingress, the state space grows quickly; bounding $|Q|$ and fixing the device set $D$ are standard ways to reflect production ingress control and avoid spurious divergence during model checking.
Appendix B AIP Message Schema Reference
To support the persistent, event-driven orchestration described in Section 7, the Agent Interaction Protocol (AIP) defines a compact set of typed message primitives that unify communication across the ConstellationClient, device agent services, and device clients. Each message carries explicit directionality (Client $\to$ Server or Server $\to$ Client), well-defined key fields, and structured reliability hooks for deterministic orchestration and safe recovery.
MsgType Dir. Key Fields Semantics Idempotent? Expected Response Reliability Hooks
[…]
HEARTBEAT C/S timestamp Liveness probe + latency sampling Yes HEARTBEAT ack (opposite dir.) HeartbeatManager jitter control
[…]
ERROR C/S error, context Protocol or execution anomaly N/A Operator / scheduler handling Rapid failure propagation
Table 5: AIP message taxonomy and reliability semantics. C=Client, S=Server. Idempotent? indicates whether duplicate delivery produces the same state (REGISTER acknowledged again, HEARTBEAT updates freshness) or is safely ignored; non-idempotent messages (e.g., COMMAND) must not be replayed without coordination.
Discussion.
The schema defines the canonical message types underlying AIP’s layered design (Section 7). REGISTER, DEVICE_INFO, and HEARTBEAT correspond to the Profile and Resilience Layers, ensuring freshness and liveness (G3, G4); TASK, COMMAND, and COMMAND_RESULTS implement deterministic execution semantics within the Execution Control Layer (G1, G5); and ERROR provides the recovery hooks required for extensibility and robustness (G6). Together, these primitives form the minimal yet expressive backbone that enables UFO3’s distributed, evolution-tolerant orchestration fabric.
Appendix C Details of NebulaBench
Table LABEL:tab:galaxy_results provides a comprehensive listing of all queries in NebulaBench, including their functional category, difficulty, the devices involved, and the observed outcomes of UFO3’s execution. Including this detailed appendix serves multiple purposes: it allows readers to (i) understand the specific nature and distribution of tasks in NebulaBench, (ii) examine UFO3’s performance at the granularity of individual queries, and (iii) enable reproducibility and comparison for future research on multi-agent orchestration and cross-device automation.
Table 6: Full NebulaBench task listing with metadata and UFO3 execution results.
ID Category Task Description Difficulty Devices Success
1 Logs Retrieve all warning and error logs from Linux servers, add them to the ’report’ sheet of the log_detailed Excel file, and send an email with the report to the operations engineer. Hard 4.0 Yes
2 Logs Search auth.log for failed SSH on all linux; create top-3 offending IPs markdown report on linux-1 Medium 3.0 Yes
3 Logs Rotate mock_app.log if >>1MB on all linux and verify shrink Medium 3.0 Yes
4 Logs Scan dmesg for OOM events on linux and return counts Medium 3.0 Yes
5 Logs Export Windows app logs and compare with Linux WARN/ERROR counts Medium 4.0 Yes
6 Logs Count how many times each systemd service was started or stopped in the last 24 hours on linux and write the results on the log_detailed excel Easy 4.0 Yes
7 Sys Detect CPU models and write best host to Notepad Easy 4.0 Yes
8 Sys Set DEMO_ENV=staging on linux-1,2 and windows; verify JSON output” Medium 3.0 No
9 Sys Ensure sandbox user exists and sudo member Easy 3.0 Yes
10 Sys Ensure backups folder 750 ops:ops on all linux Easy 3.0 Yes
11 Sys Set Windows reg flag + Linux file; confirm both paths Easy 2.0 No
12 Proc Restart cron.service and record its status Easy 2.0 Yes
13 Proc Stop cron.service on all linux Easy 3.0 Yes
14 Proc Run long_job.sh concurrently on linux 1-3 and report their running time on notepad Medium 4.0 Yes
15 Proc Get task schedule on schedule.xlsx on Windows and assign corresponding task to each linux server Hard 4.0 No
16 Proc Create hello.timer hourly on all linux, make sure it takes effect Hard 3.0 Yes
17 Data Sum today’s values from data.csv on linux-a/b; JSON output Hard 4.0 Yes
18 Data On Linux-1, Linux-2, and Linux-3, recursively scan and read their the ~/ directory on each machine for .sh files. Generate a Markdown summary table in Windows Notepad showing the file name and code summary. Medium 4.0 No
19 Data Average durations from CSV on win+linux; append metrics Medium 4.0 Yes
20 Data List /etc files on linux modified in 48h and write to the log_detailed excel” Medium 4.0 Yes
21 DevOps On all linux, run a locally available container image and ensure the container passes the health check, then close it. Hard 3.0 Yes
22 DevOps On Linux-1, clone the repository https://github.com/microsoft/UFO.git, build a Docker image named ufo:test, and push it to a shared Docker registry running on Linux-2. Then, from Linux-3, pull the same image and run a container to verify that the application starts successfully.” Hard 3.0 No
23 DevOps On all Linux, list all Docker images. Then, generate a Markdown summary on Windows Notepad showing the image list per host and highlight any differences. Hard 4.0 Yes
24 DevOps On Linux-1, clone the repository https://github.com/microsoft/UFO/ On Windows, UFO2 branch is already checked out in VSCode. Compare the UFO2 branch with Linux-1–s main branch to check if it can be merged cleanly. Report whether the merge is clean or if conflicts exist, without modifying either branch. Hard 2.0 No
25 DevOps Collect performance metrics from three Linux hosts. Based on the performance results, deploy three different microservices from their dev_path: Deploy service_1.py (lightweight) on the host with the lowest CPU load. Deploy service_2.py (medium) on the host with moderate load. Deploy service_3.py (heavy) on the host with the highest available memory. After deployment, verify that all three services are running and reachable via HTTP from each Linux node. Finally, generate a deployment report on windows notepad. Hard 4.0 Yes
26 DevOps Ensure requests repo are up to date on main branch on all linux servers Medium 3.0 Yes
27 DevOps clone https://github.com/psf/requests on all linux server on their dev path, set up the virtual environment with their agent name and install all dependencies there Hard 3.0 Yes
28 DevOps Set up a jupyer lab on linux 1 and open it with ip url on Windows browser Medium 2.0 No
29 DevOps Close all running jupyer lab linux 1. Easy 1.0 Yes
30 DevOps stop all running service on all linux Easy 3.0 Yes
31 Net Get the IP from all linux server, open the port of 8001 for linux1, 8002 for linux2, 8003 for linux3, then on each Linux server, perform ping tests to all other Linux servers on their open ports, write the results to the log_detailed excel on Windows. Hard 4.0 Yes
32 Net From Windows, check if the domain intranet.local can be resolved on both Linux-1 and Linux-2. If the domain cannot be resolved, add a temporary DNS entry for intranet.local pointing to Linux-3–s IP address in /etc/hosts on both Linux-1 and Linux-2, then verify resolution again from Windows using ping intranet.local” Medium 3.0 No
33 Net On Linux-1,check whether port 8080 is listening using ss or netstat. If it is closed, create a web service listen to this port, and, verify that it is running and listening on port 8080 from linux-2. Easy 2.0 Yes
34 Net On Linux-1, check whether http://localhost:9090/health returns a 200 OK response. If the request fails or port 9090 is closed, run a lightweight HTTP container (e.g., nginx:alpine) exposing port 9090, verify that /health returns 200 OK locally and from Linux-2, then stop and remove the container. Medium 2.0 Yes
35 Net Test scp file transfers between every pair of the four Linux nodes to ensure they can send and receive files to each other successfully. After completing all pairwise tests, write a summary report on Windows Notepad. Easy 5.0 No
36 Browsing Use the Windows browser to search for the latest stable Python release URL, download the installer, and then remotely copy and install it on Linux-a, Linux-b, and Linux-c. After installation, verify Python version on each host Hard 4.0 No
37 Browsing For all linux, get their disk usage statistics. Then, from Windows browser, search for the top 3 recommended ways to reduce high disk usage for Linux systems and document these in a report on notepad. Medium 4.0 Yes
38 Browsing Run a cpu_bench.py on all Linux. Collect results and, using Windows browser, search for recommended CPU optimization techniques for Linux servers. Create a final report on Notepad comparing benchmark results with recommended optimizations. Easy 4.0 Yes
39 Browsing Use Windows browser to download a CSV dataset of historical weather data. Copy the file to all Linux. Then, on each Linux host, compute the average temperature using a Python or shell script and save the result as weather_avg.txt Medium 4.0 No
40 Browsing From Windows browser, search for latest security CVEs for a specific Linux package. Then, on all Linux, check installed version, compare with CVE advisory, and if vulnerable, apply the patch or upgrade. Confirm patch applied successfully. Hard 4.0 Yes
41 Orchestration Complete all tasks on in schedule.xlsx on windows on linux servers, and write back the one-sentence result summary to the Result Summary column. Hard 4.0 Yes
42 Orchestration Choose lowest load host, run long_job.sh there Easy 3.0 Yes
43 Orchestration Check disk free <<10% on linux; print OK/ALERT” Easy 3.0 Yes
44 Orchestration Run cpu_bench.py on all linux hosts, and write the results to notepad on Windows Medium 4.0 Yes
45 Orchestration Create Windows hosts_summary.txt of all Linux kernels Easy 4.0 No
46 GPU check the gpu availability on the GPU node, and run the gpu_smoke.py. Summarize the results on the Notepad on Windows. Medium 2.0 Yes
47 GPU Use scp to transfer the log 1, 2, 3 from each linux machine to the gpu node and merge them to a single log_dataset.txt, then run the training.sh script. Hard 5.0 No
48 GPU ransfer the log 1, 2, 3 from each linux machine to the gpu node and merge them to a single log_dataset.txt, then run the training.sh script. Hard 4.0 Yes
49 GPU I have a distributed computation task that requires coordinating four Linux hosts. The task consists of: GPU-intensive matrix multiplication, CPU-intensive large dataset processing, Memory-intensive data aggregation. Please inspect hardware and current workload on all four hosts, to assign 1 GPU-intensive task, 2 CPU-intensive dataset processing tasks, 1 memory-intensive aggregation task to each of the host. Provide a task assignment report with hosts, hardware, assigned workload, and notes in the Notepad on Windows. Hard 5.0 Yes
50 GPU Distributedly process log files 1, 2, and 3 from each Linux machine into a format suitable for next-token prediction training. Then, merge all processed outputs into a single log_dataset.txt file on the GPU node, and execute the training.sh script. Hard 4.0 No
51 Negative Start ufo3 service on all linux Easy 3.0 No
52 Negative Deploy the service_1.py on a linux machine with 2TB memory. Medium 0.0 Yes
53 Negative Install the cuda for the linux with NVIDIA H100 GPU Medium 3.0 Yes
54 Negative Send a message to Zac on Wechat Easy 0.0 Yes
55 Negative visualize ufo3.png on linux Easy 3.0 No

### v2

@@
-UFO3: Weaving the Digital Agent Galaxy [Uncaptioned image]
+arXiv:2511.11332v2[cs.DC] 01 Mar 2026
 Abstract
@@
 1 Introduction
-The rise of intelligent agents wang2024survey marks a new era of human–computer interaction, where large language models (LLMs) naveed2025comprehensive are evolving from text-based reasoning engines qu2025tool into autonomous digital operators capable of perceiving, acting, and coordinating across tasks zhang2024large.
-Yet despite this progress, most agent frameworks remain confined within a single device or platform, be it a browser tab ning2025survey; zheng2024gpt, a desktop environment zhang2025ufo; zhang2025ufo2, or a mobile app zhang2025appagent; wang2024mobile.
+The rise of intelligent agents Wang et al.
+(2024b) marks a new era of human–computer interaction, where large language models (LLMs) Naveed et al.
+(2025) are evolving from text-based reasoning engines Qu et al.
+(2025) into autonomous digital operators capable of perceiving, acting, and coordinating across tasks Zhang et al.
+(2024).
+Yet despite this progress, most agent frameworks remain confined within a single device or platform, be it a browser tab Ning et al.
+(2025); Zheng et al.
+(2024), a desktop environment Zhang et al.
+(2025b; a), or a mobile app Zhang et al.
+(2025d); Wang et al.
+(2024a).
 This confinement sharply limits their ability to harness the rich, distributed computational ecosystem that modern users inhabit.
@@
 This gap between reasoning ability and real-world actuation leaves vast potential untapped.
-To truly advance the next frontier of automation and reasoning, agents must operate beyond the boundaries of any single device or OS, forming a coherent digital collective where Windows laptops, Linux servers, mobile devices, and edge nodes collaborate seamlessly houben2017opportunities; brudy2019cross to gather and act upon ubiquitous intelligence.
+To truly advance the next frontier of automation and reasoning, agents must operate beyond the boundaries of any single device or OS, forming a coherent digital collective where Windows laptops, Linux servers, mobile devices, and edge nodes collaborate seamlessly Houben et al.
+(2017); Brudy et al.
+(2019) to gather and act upon ubiquitous intelligence.
 Imagine a future where you could simply say: “Prepare a production-ready demo of Project X and deliver a one-page executive summary with screenshots and performance numbers.” Today, this requires tedious, error-prone coordination across devices, checking out code on a laptop, triggering GPU builds on a server, deploying to a cloud instance, recording UI interactions on a phone, and stitching results into a report.
@@
   • Declarative decomposition into a dynamic DAG (TaskConstellation).
-Natural-language or programmatic requests are decomposed by the global ConstellationAgent into a structured DAG bei2025graphs of TaskStars and TaskStarLines that encode workflow logic and dependencies.
+Natural-language or programmatic requests are decomposed by the global ConstellationAgent into a structured DAG Bei et al.
+(2025) of TaskStars and TaskStarLines that encode workflow logic and dependencies.
 This declarative structure is amenable to automated scheduling, introspection, and rewriting throughout execution.
@@
 The TaskConstellation is a living data structure.
-Intermediate outputs, transient failures, and new observations trigger controlled rewrites, adding diagnostic TaskStars, creating fallbacks, rewiring dependencies, or pruning completed nodes, so the system adapts dynamically instead of aborting on errors wu2024agentkit.
+Intermediate outputs, transient failures, and new observations trigger controlled rewrites, adding diagnostic TaskStars, creating fallbacks, rewiring dependencies, or pruning completed nodes, so the system adapts dynamically instead of aborting on errors Wu et al.
+(2024b).
   • Heterogeneous, asynchronous, and safe orchestration.
@@
 Built atop persistent WebSocket channels, we develop AIP, a protocol that provides a unified, secure, and fault-tolerant layer for agent registry, session management, task dispatch, and coordination.
-It ensures reliability under network fluctuations through automatic reconnection and retry, while exposing a lightweight, extensible interface that allows new agents to integrate seamlessly into the UFO3 ecosystem yang2025survey.
+It ensures reliability under network fluctuations through automatic reconnection and retry, while exposing a lightweight, extensible interface that allows new agents to integrate seamlessly into the UFO3 ecosystem Yang et al.
+(2025a).
   • Template-driven framework for MCP-empowered device agents.
 To democratize agent creation, UFO3 provides a lightweight development template and toolkit for rapidly building new device agents.
-Developers can declare capabilities, bind to local environments, and extend them through one or more Model Context Protocol (MCP) servers hou2025model for tool augmentation.
+Developers can declare capabilities, bind to local environments, and extend them through one or more Model Context Protocol (MCP) servers Hou et al.
+(2025) for tool augmentation.
 This modular design accelerates integration while maintaining consistency across the constellation.
@@
 Fault-injection experiments further demonstrate its robustness: UFO3 automatically retries and migrates under transient outages, gracefully degrades under partial failures, and recovers conservatively under global failures.
-In essence, UFO3 dissolves device boundaries and transforms the digital estate into a single, adaptive collaborator brudy2018investigating; marks2020multi; zhang2018towards.
+In essence, UFO3 dissolves device boundaries and transforms the digital estate into a single, adaptive collaborator Brudy et al.
+(2018); Marks and White (2020); Zhang et al.
+(2018).
 It unifies distributed devices into a cohesive digital organism, one that executes user intents safely, asynchronously, and efficiently across heterogeneous environments.
@@
 2.1 Digital Agents
-Leveraging the power of LLMs, modern digital agents have emerged as powerful interfaces bridging human intent and the complex software ecosystems people interact with daily zhang2024large.
-These agents can parse natural language requests, interpret screenshots and system state zheng2025vem; wu2025gui; zhao2025learning, decompose complex goals into subtasks huang2024understanding, and generate scripts or commands to execute tasks using a variety of tools wang2024survey.
-Execution modalities include API calls, operating-system-level commands, GUI interactions via automation or accessibility interfaces, and code generation zhang2025api; qiao2023taskweaver; wang2024large; zhang2025swe.
+Leveraging the power of LLMs, modern digital agents have emerged as powerful interfaces bridging human intent and the complex software ecosystems people interact with daily Zhang et al.
+(2024).
+These agents can parse natural language requests, interpret screenshots and system state Zheng et al.
+(2025); Wu et al.
+(2025); Zhao et al.
+(2025), decompose complex goals into subtasks Huang et al.
+(2024), and generate scripts or commands to execute tasks using a variety of tools Wang et al.
+(2024b).
+Execution modalities include API calls, operating-system-level commands, GUI interactions via automation or accessibility interfaces, and code generation Zhang et al.
+; Qiao et al.
+(2023); Wang et al.
+(2024c); Zhang et al.
+(2025e).
 Digital agents operate across a wide spectrum of platforms.
-Web-based agents can navigate browsers and search the Internet ning2025survey; zheng2024gpt, mobile agents are embedded in smartphone applications to automate mobile tasks zhang2025appagent; wang2024mobile, and desktop or laptop agents interact with local operating systems and graphical user interfaces zhang2025ufo; zhang2025ufo2.
+Web-based agents can navigate browsers and search the Internet Ning et al.
+(2025); Zheng et al.
+(2024), mobile agents are embedded in smartphone applications to automate mobile tasks Zhang et al.
+(2025d); Wang et al.
+(2024a), and desktop or laptop agents interact with local operating systems and graphical user interfaces Zhang et al.
+(2025b; a).
 Across these platforms, agents observe system state, reason about tasks, and execute actions, enabling applications such as automated customer support flows, desktop productivity macros, and repetitive workflow automation, essentially acting as intelligent digital assistants.
 Despite their versatility and growing adoption, existing digital agents are fundamentally limited by their confinement to a single device or environment.
-In practice, many modern workflows are no longer confined to a single endpoint brudy2018investigating: users frequently interact with a combination of desktops, mobile devices, cloud services, and specialized hardware, all as part of a single task xu2021cross; chen2020multi.
+In practice, many modern workflows are no longer confined to a single endpoint Brudy et al.
+(2018): users frequently interact with a combination of desktops, mobile devices, cloud services, and specialized hardware, all as part of a single task Xu et al.
+(2021); Chen and Koufaris (2020).
 Examples include running data analysis on a GPU cluster, collecting results on a personal laptop, and generating visual summaries on a tablet, or coordinating cross-platform deployments that touch both local workstations and cloud infrastructure.
@@
   • Fragmented personalization and context.
-User-specific preferences, personalization, and contextual knowledge captured on one device are difficult to transfer or leverage on another cemri2025multi, preventing a coherent multi-device experience.
+User-specific preferences, personalization, and contextual knowledge captured on one device are difficult to transfer or leverage on another Cemri et al.
+(2025), preventing a coherent multi-device experience.
   • Manual coordination overhead.
@@
 2.2 Cross-Device Agent: A New Paradigm
-To overcome the limitations of single-device frameworks, we introduce the concept of cross-device intelligence, in which digital agents collaborate seamlessly across multiple heterogeneous endpoints brudy2019cross.
+To overcome the limitations of single-device frameworks, we introduce the concept of cross-device intelligence, in which digital agents collaborate seamlessly across multiple heterogeneous endpoints Brudy et al.
+(2019).
 In this metaphorical digital galaxy, each device, desktop, mobile, cloud service, or specialized hardware, acts as a star, and coordinated tasks form a constellation that collectively fulfills complex user requests.
-Unlike traditional agents confined to a single host, these cross-device agents can reason about device capabilities, orchestrate subtasks, and execute actions as part of a unified, intelligent ecosystem cheninternet
+Unlike traditional agents confined to a single host, these cross-device agents can reason about device capabilities, orchestrate subtasks, and execute actions as part of a unified, intelligent ecosystem Chen et al.
 Cross-device intelligence enables a qualitatively new user experience.
@@
 It overcomes the constraints of individual devices by assigning tasks to the endpoints with the necessary capabilities and resources.
-It allows personalization, context, and user preferences captured on one device to propagate across others, enabling a coherent multi-device experience giusti2025federation.
+It allows personalization, context, and user preferences captured on one device to propagate across others, enabling a coherent multi-device experience Giusti et al.
+(2025).
 And it automates orchestration of complex workflows, removing reliance on brittle glue code or labor-intensive manual integration.
@@
 However, synchronous control loops, cross-device agent orchestration introduces a fundamentally different class of systems challenges.
-In a distributed, heterogeneous environment, agents must collaborate across device, network, and platform boundaries, each with distinct runtime contexts and execution semantics tran2025multi.
+In a distributed, heterogeneous environment, agents must collaborate across device, network, and platform boundaries, each with distinct runtime contexts and execution semantics Tran et al.
+(2025).
 Building a reliable and efficient orchestration layer under such conditions requires addressing three core challenges.
@@
 Unlike linear single-agent plans, task execution must accommodate partial completions, delayed feedback, and dynamic dependency resolution.
-The orchestrator must therefore reason about concurrency, detect when subtasks can safely proceed in parallel, and continuously adapt its scheduling plan based on evolving runtime states and network latencies yu2025dyntaskmas.
+The orchestrator must therefore reason about concurrency, detect when subtasks can safely proceed in parallel, and continuously adapt its scheduling plan based on evolving runtime states and network latencies Yu et al.
+(2025).
 Failing to do so may lead to wasted computation or stalled progress.
@@
 Achieving coherent coordination among them requires standardized mechanisms for registration, capability advertisement, task dispatch, and result collection.
-These operations must occur through persistent, low-latency channels that can tolerate temporary disconnections and guarantee consistent task states cheninternet.
+These operations must occur through persistent, low-latency channels that can tolerate temporary disconnections and guarantee consistent task states Chen et al.
+.
 In practice, ad-hoc HTTP calls or ephemeral connections are insufficient; a structured communication substrate is essential to sustain long-lived agent interactions.
@@
 Device agents differ in operating systems, execution environments, and available toolchains.
-The architecture should therefore allow rapid development, deployment, and integration of new device agents while ensuring consistent semantics in task execution and error propagation wu2024autogen.
-This requires a flexible interface model and a unified protocol to abstract platform-specific complexity, enabling scalable and evolvable orchestration across an open agent federation yang2025survey.
+The architecture should therefore allow rapid development, deployment, and integration of new device agents while ensuring consistent semantics in task execution and error propagation Wu et al.
+(2024a).
+This requires a flexible interface model and a unified protocol to abstract platform-specific complexity, enabling scalable and evolvable orchestration across an open agent federation Yang et al.
+(2025a).
 These challenges are not merely engineering details, as they delineate the boundary between ad-hoc agent scripts and a principled, distributed orchestration system.
@@
 4.4 Runtime Dynamism and Adaptivity
-Unlike static DAG schedulers Islam2012Oozie; di2017nextflow; argo, UFO3 treats TaskConstellations as mutable objects.
+Unlike static DAG schedulers M.
+Islam, A.
+K.
+Huang, M.
+Battisha, M.
+Chiang, S.
+Srinivasan, C.
+Peters, A.
+Neumann, and A.
+Abdelnur (2012); P.
+Di Tommaso, M.
+Chatzou, E.
+W.
+Floden, P.
+P.
+Barja, E.
+Palumbo, and C.
+Notredame (2017); 2, UFO3 treats TaskConstellations as mutable objects.
 Tasks and dependency edges can be inserted, removed, or modified at runtime.
@@
 As illustrated in Figure 4, ConstellationAgent acts as both a planner and a replanner: it first synthesizes a TaskConstellation from user instructions, then incrementally refines this constellation as feedback arrives from distributed agents.
-Internally, ConstellationAgent is implemented as an LLM-driven ReAct agent yao2022react governed by a finite-state machine (FSM) schneider2005state.
+Internally, ConstellationAgent is implemented as an LLM-driven ReAct agent Yao et al.
+(2022) governed by a finite-state machine (FSM) Schneider (2005).
 This FSM alternates between two complementary operating modes, namely creation and editing, forming a closed-loop control cycle that continuously updates the global execution graph in response to runtime conditions.
@@
 Creation Mode.
-In the creation phase, ConstellationAgent receives three primary inputs: (i) a user-issued goal, expressed in natural or structured language; (ii) the AgentProfile registry, describing each available device agent’s capabilities, environment, and metadata; and (iii) demonstration examples to support in-context learning (ICL) dong2022survey; jiang2024xpert.
+In the creation phase, ConstellationAgent receives three primary inputs: (i) a user-issued goal, expressed in natural or structured language; (ii) the AgentProfile registry, describing each available device agent’s capabilities, environment, and metadata; and (iii) demonstration examples to support in-context learning (ICL) Dong et al.
+(2022); Jiang et al.
+(2024).
 Leveraging the LLM’s semantic reasoning capabilities, ConstellationAgent decomposes the user goal into a structured execution graph, the TaskConstellation.
@@
 The TaskConstellation is generated in a structured, machine-readable JSON format.
-Beyond the raw graph, ConstellationAgent produces a detailed reasoning trace: it first formulates an Observation of the input and AgentProfile, followed by a structured Thought capturing its analysis and decision-making rationale wei2022chain; ding2024everything.
+Beyond the raw graph, ConstellationAgent produces a detailed reasoning trace: it first formulates an Observation of the input and AgentProfile, followed by a structured Thought capturing its analysis and decision-making rationale Wei et al.
+(2022); Ding et al.
+(2024).
 It then outputs the next State in the control loop and the overall Result for user-request or the generated TaskConstellation.
@@
 build_constellation Batch-create tasks and dependencies from structured input Configuration dictionary, clear flag Built TaskConstellation
-To operationalize dynamic graph construction, ConstellationAgent interacts with a lightweight Constellation MCP Server hou2025model, a modular component exposing a standardized set of task and dependency management primitives.
+To operationalize dynamic graph construction, ConstellationAgent interacts with a lightweight Constellation MCP Server Hou et al.
+(2025), a modular component exposing a standardized set of task and dependency management primitives.
 This server serves as the structured manipulation layer that bridges LLM-level reasoning and concrete execution state.
@@
 Figure 6: The Constellation Orchestrator bridges TaskConstellation and execution, enabling asynchronous, adaptive task orchestration across devices.
-With the ConstellationAgent governs the reasoning and evolution of a TaskConstellation, the Constellation Orchestrator brings that plan to life by executing, monitoring, and adapting interdependent tasks across heterogeneous devices li2022dag, as shown in Figure 6.
+With the ConstellationAgent governs the reasoning and evolution of a TaskConstellation, the Constellation Orchestrator brings that plan to life by executing, monitoring, and adapting interdependent tasks across heterogeneous devices Li et al.
+(2022), as shown in Figure 6.
 Conceptually, it transforms a static DAG into a living execution fabric, where tasks evolve concurrently, react to runtime signals, and adapt to new decisions generated by the reasoning agent.
@@
 The orchestration model described in Section 6 requires a communication substrate that remains correct under continuous DAG evolution, dynamic agent participation, and fine-grained event propagation.
-Legacy HTTP-based coordination approaches (e.g., A2A duan2025agent, ACP ehtesham2025survey) assume short-lived, stateless interactions, incurring handshake overhead, stale capability views, and fragile recovery when partial failures occur mid-task.
+Legacy HTTP-based coordination approaches (e.g., A2A Duan and Lu (2025), ACP Ehtesham et al.
+(2025)) assume short-lived, stateless interactions, incurring handshake overhead, stale capability views, and fragile recovery when partial failures occur mid-task.
 These assumptions make them unsuitable for the continuously evolving workflows and long-running reasoning loops characteristic of UFO3.
@@
 Following the system implementation, we evaluate the UFO3 framework to understand its effectiveness, robustness, and coordination efficiency in realistic, heterogeneous environments.
-Existing agent benchmarks predominantly focus on single-device or single-OS settings (e.g., text-based API workflows or GUI automation) mu2025gui, which fail to capture the cross-device, cross-platform orchestration that UFO3 is designed for.
+Existing agent benchmarks predominantly focus on single-device or single-OS settings (e.g., text-based API workflows or GUI automation) Mu et al.
+(2025), which fail to capture the cross-device, cross-platform orchestration that UFO3 is designed for.
 Moreover, to the best of our knowledge, there exists no prior agent system capable of orchestrating multi-device tasks that span both Linux and Windows environments with unified control and shared context.
@@
 Agents communicate via a simulated local network with 1–10 ms latency and a wide-area link (50–100 ms latency) for the GPU node using the AIP.
-All components are implemented in Python 3.10 and powered by the same large language model (GPT-5-Chat-20251003) openai2025gpt5systemcard for both the ConstellationAgent and device agents.
+All components are implemented in Python 3.10 and powered by the same large language model (GPT-5-Chat-20251003) OpenAI (2025b) for both the ConstellationAgent and device agents.
 The orchestrator and controller run on a dedicated management node that coordinates agent discovery, task scheduling, and monitoring.
@@
 Table 3: Average modifications per edit and per request by type.
-Added
-Tasks
-Removed
-Tasks
-Modified
-Tasks
-Added
-Dependencies
-Removed
-Dependencies
-Modified
-Dependencies
-Total
+Added Tasks Removed Tasks Modified Tasks Added Dependencies Removed Dependencies Modified Dependencies Total
 Per Edit 0.05 0.00 0.99 0.04 0.01 0.00 1.09
@@
 Table 4: Parallelism characteristics of task constellations by difficulty.
-Difficulty
-Max
-Parallel Width
-Critical Path
-Length (L)
-Total Exec.
-Time (W)
-Parallelism
-Ratio (P = W/L)
+Difficulty Max Parallel Width Critical Path Length (L) Total Exec. Time (W) Parallelism Ratio (P = W/L)
 Easy 3.53 144.10 315.28 1.86
@@
 Moreover, UFO3 orchestrates at the protocol level rather than binding to specific implementations.
-Any agent compatible with AIP, whether it is a coding agent yang2024swe, a deep-research agent huang2025deep, or a data-analysis module zhang2025allhands, can participate in the orchestration.
+Any agent compatible with AIP, whether it is a coding agent Yang et al.
+(2024), a deep-research agent Huang et al.
+(2025), or a data-analysis module Zhang et al.
+(2025c), can participate in the orchestration.
 This design opens the door for a broader agent ecosystem, where specialized agents cooperate under a unified orchestration fabric.
@@
 However, the system’s overall capability remains constrained by the competence of individual device agents.
-Since execution is distributed, a single agent’s failure, such as incomplete subtask execution or incorrect result reporting, can cascade through the constellation and lead to global task failure cemri2025multi.
+Since execution is distributed, a single agent’s failure, such as incomplete subtask execution or incorrect result reporting, can cascade through the constellation and lead to global task failure Cemri et al.
+(2025).
 This “weakest-link effect” underscores the need for more reliable and self-verifying device agents.
@@
 Many cross-device tasks, however, require file transfer or shared context, for example, a Windows agent generating a dataset for a Linux node to process.
-To support such workflows, we plan to extend AIP with a persistent, shared “agent memory” zhang2025survey, effectively a distributed workspace where agents can store and retrieve intermediate results.
+To support such workflows, we plan to extend AIP with a persistent, shared “agent memory” Zhang et al.
+(2025f), effectively a distributed workspace where agents can store and retrieve intermediate results.
 This enhancement would unify context across devices, simplify file sharing, and enable richer collaborative behavior across heterogeneous agents.
@@
 12.1 Digital Device Agents
-The emergence of large language models, particularly multimodal variants, has enabled a new class of intelligent agents capable of operating directly on digital devices zhang2024large.
+The emergence of large language models, particularly multimodal variants, has enabled a new class of intelligent agents capable of operating directly on digital devices Zhang et al.
+(2024).
 These agents extend beyond web and mobile environments to desktops, and are now expanding into domains such as automotive and embedded systems.
-Early systems like SeeAct zheng2024gpt, Mobile-Agent wang2024mobile, and UFO zhang2025ufo pioneered this direction by leveraging GPT-4V yang2023dawn to perceive graphical user interfaces and perform human-like interactions across different platforms.
-Subsequent research has evolved along two main trajectories, namely (i) system-level integration, enhancing robustness through tighter coupling with native OS and API layers zhang2025api; zhang2025ufo2; and (ii) model-level adaptation, fine-tuning LLMs with large-scale interaction data and reinforcement learning to improve grounding and autonomy in digital environments luo2025gui; zheng2025vem; wang2025ui.
+Early systems like SeeAct Zheng et al.
+(2024), Mobile-Agent Wang et al.
+(2024a), and UFO Zhang et al.
+(2025b) pioneered this direction by leveraging GPT-4V Yang et al.
+(2023) to perceive graphical user interfaces and perform human-like interactions across different platforms.
+Subsequent research has evolved along two main trajectories, namely (i) system-level integration, enhancing robustness through tighter coupling with native OS and API layers Zhang et al.
+; Zhang et al.
+(2025a); and (ii) model-level adaptation, fine-tuning LLMs with large-scale interaction data and reinforcement learning to improve grounding and autonomy in digital environments Luo et al.
+(2025); Zheng et al.
+(2025); Wang et al.
+(2025).
 Industry adoption has accelerated rapidly.
-Anthropic’s Claude (Computer Use) anthropic2024 relies purely on screenshot-based perception, while OpenAI’s Operator cua2025 demonstrates strong multimodal reasoning and robust desktop automation.
+Anthropic’s Claude (Computer Use) Anthropic (2024) relies purely on screenshot-based perception, while OpenAI’s Operator OpenAI (2025a) demonstrates strong multimodal reasoning and robust desktop automation.
 These advances collectively signal the formation of a new computing paradigm where LLMs serve as universal control interfaces for heterogeneous software systems.
@@
 12.2 Multi-Agent Orchestration
-Recent years have witnessed growing interest in LLM-based multi-agent systems, where coordination among heterogeneous agents becomes central to achieving emergent intelligence guo2024large.
-Designing robust orchestration protocols, covering role assignment, communication, collaboration, and context sharing, has been identified as a key challenge bhatt2025should; dang2025multi; kong2025survey.
-The Internet of Agents (IoA) cheninternet represents one of the earliest structured attempts in this direction.
+Recent years have witnessed growing interest in LLM-based multi-agent systems, where coordination among heterogeneous agents becomes central to achieving emergent intelligence Guo et al.
+(2024).
+Designing robust orchestration protocols, covering role assignment, communication, collaboration, and context sharing, has been identified as a key challenge Bhatt et al.
+(2025); Dang et al.
+(2025); Kong et al.
+(2025).
+The Internet of Agents (IoA) Chen et al.
+represents one of the earliest structured attempts in this direction.
 It introduces an integration protocol and an instant-messaging-style communication layer, enabling flexible, scalable multi-agent collaboration.
-IoA emphasizes dynamic teaming and conversation flow control, allowing agents to self-organize and cooperate fluidly in an Internet-like environment yang2025agentic.
-Building upon this idea, the Federation of Agents (FoA) giusti2025federation further advances distributed orchestration with Versioned Capability Vectors (VCVs) as machine-readable profiles that encode agent capabilities, costs, and constraints.
+IoA emphasizes dynamic teaming and conversation flow control, allowing agents to self-organize and cooperate fluidly in an Internet-like environment Yang et al.
+(2025b).
+Building upon this idea, the Federation of Agents (FoA) Giusti et al.
+(2025) further advances distributed orchestration with Versioned Capability Vectors (VCVs) as machine-readable profiles that encode agent capabilities, costs, and constraints.
 These semantic embeddings make capabilities discoverable and composable, enabling dynamic, capability-driven coordination at scale.